### v1

# Complexity reduction of physical models: an equation-free approach by means of scaling

Abstract
The description of complex physical phenomena often involves sophisticated models that rely on a large number of parameters, with many dimensions and scales. One practical way to simplify that kind of models is to discard some of the parameters, or terms of underlying equations, thus giving rise to reduced models. Here, we propose a general approach to obtaining such reduced models. The method is independent of the model in use, i.e., equation-free, depends only on the interplay between the scales and dimensions involved in the description of the phenomena, and controls over-parametrization. It also quantifies conditions for asymptotic models by providing explicitly computable thresholds on values of parameters that allow for reducing complexity of a model, while preserving essential predictive properties. Although our focus is on complexity reduction, this approach may also help with calibration by mitigating the risks of over-parameterization and instability in parameter estimation. The benefits of this approach are discussed in the context of the classical projectile model.
journal: Proc. R. Soc. A
1 Introduction
When dealing with complex models, it is often useful to consider an asymptotic analysis with the aim of simplifying the model under consideration. Such an approach identifies a parameter $\varepsilon>0$ and investigates the limit for $\varepsilon\to 0$ of the solutions to the given equations. When $\varepsilon$ is small, the solutions to the model considered in first place can be accurately reproduced by the solutions of the same equations with $\varepsilon$ formally set to 00, i.e., with some terms neglected. However, in general, it is hard to understand how small such an $\varepsilon$ must be to consider it “small enough” and, then, to allow the use of solutions to the simplified model.
In particular, for models involving very different scales or over-parametrization, reducing complexity and applying appropriate scaling becomes crucial for preventing instability in calibration [1, 2, 3]. This turns to be even more critical when feeding synthetic data from physical models into machine learning models [4], as imbalances or miscalibrations in the data can significantly affect model performance [5]. Proper scaling ensures that synthetic data reflects realistic parameter distributions, helping to mitigate overfitting and instability in parameter estimation [1]. Ultimately, scaling improves calibration, stabilizes predictions and may lead to more reliable machine learning model training, even when using synthetic data.
Our approach is based on already introduced methodologies, that we called Optimal Scaling (OS) [6] and Optimal Scaling with Constraints (OSC) [7]. Similarly to traditional nondimensionalization approaches, those allow recasting dimensional equations into unitless form. However, in contrast to traditional techniques, they do not require a problem-dependent insight into the phenomenon under consideration. In addition, they ensure a strong reduction of ranges of magnitudes assumed by the equations’ coefficients, provide accurate estimation of characteristic features of considered systems and allow reducing round-off and numerical errors.
Such approaches were applied in [6, 7] to the prospective Population Balance Model [8] for Latex Particles Morphology Formation (LPMF PBM) [9, 10]. First, in [6], we obtained dimensionless scaled equations with minimized variation of the relevant quantities. Such a feature was beneficial in diminishing numerical errors, and avoiding unphysical behavior of the computed solution. Second, in [7], a quantitative criterion for locating regions of slow and fast aggregation was introduced to derive a family of dimensionless LPMF PBM of reduced complexity. Moreover, in [6], the Optimal Scaling was also applied to liquid crystal and quantum models, helping to decrease the range of magnitudes assumed by the involved quantities, without compromising on accuracy of characteristic features of considered systems [11, 12].
Here, we present an advanced variant of the Optimal Scaling approaches which inherits benefits of its predecessors and, in addition, takes control of over-parametrization. We refer to this method as Generalized Optimal Scaling (GOS). The proposed methodology can provide a quantitative criterion for identifying those values of physical parameters that allow discarding a few terms in the considered equations without significantly modifying their solutions. This technique does not depend on the model, and as such its application to any specific model will produce results which exhibit a certain dependency on the model, that needs to be investigated numerically. In order to concretely illustrate our derivations and findings, we consider the classical projectile model.
The paper is structured as follows. Section 2 clarifies the employed notations and summarizes the projectile model we use in this work as a case study. Section 3 outlines the main achievements of the proposed scaling methods [6, 7] and puts the OSC method, previously formulated in the context of the LPMF PBM, into a general framework. Section 4 improves further the generalized OSC by extending its functionalities with ability to reduce to its minimum the number of parameters employed in a model. A detailed formulation of the resulting GOS method is presented in Table 1 in comparison with the original OSC method. In Section 5, we test the Optimal Scaling methodologies on the projectile model. Section 6 provides conclusions and discusses the capabilities of presented techniques.
2 Prerequisites
2.1 Employed Notation
Through all this work, the following notation is employed.
  • We denote any scalar quantity $\tilde{q}$ as
    $\tilde{q}\coloneqq q\,[\tilde{q}],$ (1)
    being $[\tilde{q}]$ the dimensions of $\tilde{q}$, i.e., its units of measure, and $q$ the numerical value associated with $\tilde{q}$ in the chosen units $[\tilde{q}]$.
  • The symbol $\ll$ is used when a strictly positive number $a$ is at least an order of magnitude smaller than another strictly positive number $b$, i.e., $a\ll b\Leftrightarrow a/b\leq 10^{-1}$.
  • The symbol $\approx$ is used when the two strictly positive numbers $a$ and $b$ have similar magnitudes, i.e., $a\approx b\Leftrightarrow 10^{-1}<a/b<10$.
  • The superscript $T$ means “transpose”, e.g., the matrix $\textbf{M}^{T}$ is the transpose of M.
  • The acronyms OS, OSC and GOS refer to Optimal Scaling, Optimal Scaling with Constraints and Generalized Optimal Scaling, respectively.
2.2 The classical projectile model: scaling strategies
We choose the classical projectile model as an application for the scaling approaches considered in this work. First, we wish to discuss some possible strategies for scaling this model in the context of the Optimal Scaling to highlight the ideas behind the scaling methods under study. The rigorous scaling of the model is presented in detail in Section 5.
A ball is thrown vertically upward from a certain height above the surface of the Earth. The height $\tilde{y}(\tilde{t})$ reached at time $\tilde{t}$ is measured in meters, while the elapsed time $\tilde{t}$ in seconds, i.e., $[\tilde{y}]=\mathrm{m}$, $[\tilde{t}]=\mathrm{s}$. As presented in [13], the variable $\tilde{y}$ satisfies the Ordinary Differential Equation
$\frac{d^{2}\tilde{y}}{d\tilde{t}^{2}}=-\frac{\tilde{g}\,\tilde{R}^{2}}{(\tilde{y}+\tilde{R})^{2}},\quad\forall\tilde{t}\geq\tilde{t}_{0}\coloneqq 0\,\mathrm{s},\quad\mbox{with}\quad\tilde{y}(\tilde{t}_{0})=\tilde{h}_{0},\quad\frac{d\tilde{y}}{d\tilde{t}}(\tilde{t}_{0})=\tilde{v}_{0},$ (2)
being $\tilde{g}$ the gravitational acceleration, $\tilde{R}$ the radius of the Earth, $\tilde{h}_{0}$ and $\tilde{v}_{0}$ the initial height and velocity, respectively. As indicated in [14], the dimensionless counterpart of (2) is obtained by the change of variables
$\tau\coloneqq\tilde{t}/\tilde{\theta}_{1},\quad\varphi\coloneqq\tilde{y}/\tilde{\theta}_{2},\quad[\tilde{\theta}_{1}]=[\tilde{t}],\quad[\tilde{\theta}_{2}]=[\tilde{y}],\quad\theta_{1},\theta_{2}\in(0,\infty),$ (3)
where $\tilde{\theta}_{1}$ and $\tilde{\theta}_{2}$ are characteristic constants of the considered model (2). The transformation (3) leads to
$\frac{d^{2}\varphi}{d\tau^{2}}=-\frac{\lambda_{1}}{(1+\lambda_{2}\,\varphi)^{2}},\quad\forall\tau\geq 0,\quad\mbox{with}\quad\varphi(0)=\lambda_{3},\quad\frac{d\varphi}{d\tau}(0)=\lambda_{4}.$ (4)
Here, the dimensionless coefficients $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}\in(0,\infty)$ are defined as
$\lambda_{1}\coloneqq\tilde{p}_{1}\,\tilde{\theta}_{1}^{2}\,\tilde{\theta}_{2}^{-1},\quad\lambda_{2}\coloneqq\tilde{p}_{2}^{-1}\,\tilde{\theta}_{2},\quad\lambda_{3}\coloneqq\tilde{p}_{3}\,\tilde{\theta}_{2}^{-1},\quad\lambda_{4}\coloneqq\tilde{p}_{4}\,\tilde{\theta}_{1}\,\tilde{\theta}_{2}^{-1},$ (5)
with
$\tilde{p}_{1}\coloneqq\tilde{g},\quad\tilde{p}_{2}\coloneqq\tilde{R},\quad\tilde{p}_{3}\coloneqq\tilde{h}_{0},\quad\tilde{p}_{4}\coloneqq\tilde{v}_{0}.$ (6)
The scaling (3) results in $N_{\mathsf{l}}=4$ dimensionless coefficients $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}$ (5), computed as functions of $N_{\mathsf{x}}=2$ characteristic constants $\tilde{\theta}_{1},\tilde{\theta}_{2}$ and $N_{\mathsf{p}}=4$ predefined physical parameters $\tilde{p}_{1},\tilde{p}_{2},\tilde{p}_{3},\tilde{p}_{4}$ (6). One can now optimize some of dimensionless coefficients, i.e., $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}$, as functions of $\tilde{\theta}_{1},\tilde{\theta}_{2}$, which is the path followed by OS and OSC.
Alternatively, one can reduce the number of physical parameters to just two, namely
$\pi_{1}=\tilde{g}^{-1/2}\,\tilde{R}^{-1/2}\,\tilde{v}_{0},\quad\pi_{2}=\tilde{R}^{-1}\,\tilde{h}_{0}.$ (7)
Now it is possible to optimize, say, the coefficients $\lambda_{2},\lambda_{3},\lambda_{4}$ as functions of $\pi_{1},\pi_{2}$ (instead of $\tilde{\theta}_{1},\tilde{\theta}_{2}$, as suggested in [6, 7]). If one aims to have $\lambda_{2},\lambda_{3},\lambda_{4}$ as close as possible to $1$ (the approach taken by OS), it follows:
$\lambda_{2}=\pi_{2}^{1/2},\quad\lambda_{3}=\pi_{2}^{1/2},\quad\lambda_{4}=1.$ (8)
Then, we will necessarily have that $\lambda_{1}=\pi_{1}^{-2}\,\pi_{2}^{1/2}$. Further, introducing the condition $\lambda_{1}\ll 1$ as considered by OSC (which in concrete numerical computations we take to be $\lambda_{1}\leq 10^{-1}$) provides a necessary relationship between $\pi_{1}$ and $\pi_{2}$, a threshold below which the $\lambda_{1}$ can be discarded and results in a reduced model.
It should be noted that the analysis of the $\lambda_{i}$, $i=1,2,3,4$, as functions of $\pi_{1},\pi_{2}$, is based on just dimensional considerations, and is thus equation-free and model independent. Indeed, one needs afterwards a specific study, based at very least on numerical simulations, in order to identify the precise meaning of the approximation of the solutions of (4) by the reduced model in which $\lambda_{1}$ is set to be zero.
We return to the projectile model in Section 5 in the context of the novel scaling method to be introduced below.
3 Optimal Scaling: Overview
We consider models that can be formulated in terms of
[…]
where $\tilde{F}$ is a function (with dimensions $[\tilde{F}]$) of unknown variables and independent quantities $\tilde{x}_{1},\dots,\tilde{x}_{N_{\mathsf{x}}}$ and physical parameters $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$, whose numerical values
$F,x_{1},\dots,x_{N_{\mathsf{x}}}\in\mathbb{C},\quad p_{1},\dots,p_{N_{\mathsf{p}}}\in(0,\infty).$ (10)
Aiming to rewrite (9) in unitless form, we define the dimensionless counterparts of $\tilde{x}_{1},\dots,\tilde{x}_{N_{\mathsf{x}}}$ as:
$\xi_{1}\coloneqq\tilde{x}_{1}/\tilde{\theta}_{1},\dots,\xi_{N_{\mathsf{x}}}\coloneqq\tilde{x}_{N_{\mathsf{x}}}/\tilde{\theta}_{N_{\mathsf{x}}},\quad\mbox{with}\quad[\tilde{\theta}_{j}]=[\tilde{x}_{j}],\quad\theta_{j}\in(0,\infty),\quad\forall j=1,\dots,N_{\mathsf{x}},$ (11)
being $\tilde{\theta}_{1},\dots,\tilde{\theta}_{N_{\mathsf{x}}}$ the characteristic constants of the dimensional model (9). By applying (11), it is possible to rewrite (9) in unitless form
$f\left(\xi_{1},\dots,\xi_{N_{\mathsf{x}}},\lambda_{1},\dots,\lambda_{N_{\mathsf{l}}}\right)=0,\quad f:\mathbb{C}^{N_{\mathsf{x}}}\times(0,\infty)^{N_{\mathsf{l}}}\to\mathbb{C},$ (12)
with $\lambda_{1},\dots,\lambda_{N_{\mathsf{l}}}\in(0,\infty)$ being the dimensionless coefficients in (12). The coefficients $\lambda_{1},\dots,\lambda_{N_{\mathsf{l}}}$ in (12) depend only on the characteristic constants $\tilde{\theta}_{1},\dots,\tilde{\theta}_{N_{\mathsf{x}}}$ and physical parameters $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$, i.e.,
$\lambda_{i}=\lambda_{i}(\tilde{\theta}_{1},\dots,\tilde{\theta}_{N_{\mathsf{x}}},\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}})\in(0,\infty),\quad\forall i=1,\dots,N_{\mathsf{l}}.$ (13)
Given a fixed choice of physical parameters $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$, the numerical values $\theta_{1},\dots,\theta_{N_{\mathsf{x}}}\in(0,\infty)$ of the characteristic constants $\tilde{\theta}_{1},\dots,\tilde{\theta}_{N_{\mathsf{x}}}$ should be determined to complete the definition of the dimensionless model (12).
In the literature [13, 14], it is often suggested to set $N_{\mathsf{x}}$ out of the $N_{\mathsf{l}}$ coefficients $\lambda_{1},\dots,\lambda_{N_{\mathsf{l}}}$ equal to $1$. Then, one aims to solve the resulting system of equations with respect to the values $\theta_{1},\dots,\theta_{N_{\mathsf{x}}}$ of the characteristic constants. This approach presents some deficiencies, such as, e.g., at most $N_{\mathsf{x}}$ out of $N_{\mathsf{l}}$ independent quantities can be enforced to the value of $1$, with no control on the rest of the dimensionless coefficients. Furthermore, the selection of coefficients enforced to the value of $1$ depends on the particular problem with no general guidelines given so far.
In this paper, we address these issues by proposing techniques for selection of numerical values $\theta_{1},\dots,\theta_{N_{\mathsf{x}}}\in(0,\infty)$ of the characteristic constants $\tilde{\theta}_{1},\dots,\tilde{\theta}_{N_{\mathsf{x}}}$ to scale the dimensional model (9) effectively, according to the objectives set for the specific scaling method (see Table 1 from Section 4). We remark that such procedures do not require a physical insight into the considered phenomenon and corresponding equations. The main features and achievements of presented methodologies are summarized in what follows.
3.1 Optimal Scaling
Optimal Scaling (OS) proposed in [6] enables such a rational choice of $\theta_{1},\dots,\theta_{N_{\mathsf{x}}}\in(0,\infty)$, that ensures the deviation of magnitudes of $\lambda_{1},\dots,\lambda_{N_{\mathsf{l}}}$ (13) from unity to be minimal for any fixed choice of physical parameters $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$. This is advantageous for numerical computations as the floating-point numbers have the highest density in the interval $(0,1)$ and their density decreases when moving further away from this interval. So, it is always convenient to manipulate numerically the quantities with magnitude $\approx 1$, in order to avoid severe round-off errors and to improve the conditioning of the problem at hand.
In [6], such a methodology was first applied to liquid crystal and quantum models. The reported data demonstrate that OS ensures a strong reduction of the range of orders of magnitude assumed by involved quantities. Despite such a strong transformation, OS allows for an accurate estimation of characteristic features of considered systems. As an example, we showed that OS correctly describes the large-body limit of Landau-de Gennes model for liquid crystals [11]. Moreover, for Schrödinger Equation [12], we proved that the characteristic constants, provided by OS, follow the same trend as the often used atomic units.
Further, it was demonstrated that OS was also beneficial for the case of the Population Balance Model [8] for Latex Particles Morphology Formation [9, 10]. When compared with other scaling approaches [13, 14], it assured the smallest variation $\mathop{\mathrm{max}}_{i=1,\dots,N_{\mathsf{l}}}\lambda_{i}/\mathop{\mathrm{min}}_{i=1,\dots,N_{\mathsf{l}}}\lambda_{i}$ of coefficients (13) and allowed reducing round-off and numerical errors, as discussed in [6].
3.1.1 Formulation
Given any fixed choice of physical parameters $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$, OS proposed in [6] finds $\vec{\theta}=\vec{\theta}_{\mathrm{opt}}$ in such a way that the deviation of magnitudes of $\lambda_{1},\dots,\lambda_{N_{\mathsf{l}}}$ (13) from unity is minimal, i.e.,
$\vec{\theta}_{\mathrm{opt}}\coloneqq\mathop{\mathrm{argmin}}_{\vec{\theta}\in(0,\infty)^{N_{\mathsf{x}}}}C(\vec{\theta}),\quad C(\vec{\theta})\coloneqq\sum_{i=1}^{N_{\mathsf{l}}}\left(\log_{10}(\lambda_{i}(\tilde{\theta}_{1},\dots,\tilde{\theta}_{N_{\mathsf{x}}},\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}))-\Theta_{i}\right)^{2},$
$\mbox{where}\quad\vec{\theta}\coloneqq(\theta_{1},\dots,\theta_{N_{\mathsf{x}}})\in(0,\infty)^{N_{\mathsf{x}}}\quad\mbox{and the constants}\quad\Theta_{1}=\cdots=\Theta_{N_{\mathsf{l}}}=0$ (14)
hold the targeted orders of magnitude, i.e., $\Theta_{i}=\log_{10}(1)=0$, $\forall i=1,\dots,N_{\mathsf{l}}$.
With $U=\{1,\dots,N_{\mathsf{l}}\}$, the function $C$ (14) is equivalent to $C_{U}$ in Theorem B.1 (Appendix B). Then, given Property A.1 (Appendix A) and any fixed choice of $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$, the points $\vec{\theta}_{\mathrm{opt}}$ of global minimum of $C\equiv C_{U}$ can be computed by means of (B.2), as follows from Theorem B.1 with $U=\{1,\dots,N_{\mathsf{l}}\}$. In this way, OS provides scaled coefficients $\lambda_{1},\dots,\lambda_{N_{\mathsf{l}}}$ (13) with magnitudes as close as possible to the targeted value $1$. In other words, OS attempts to make
$\lambda_{i}(\tilde{\theta}_{1},\dots,\tilde{\theta}_{N_{\mathsf{x}}},\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}})\approx 1,\quad\mbox{when}\quad\tilde{\theta}_{j}=\vec{\theta}_{\mathrm{opt},j}\,[\tilde{\theta}_{j}],\quad\forall i,j,$ (15)
with $\vec{\theta}_{\mathrm{opt},j}$ being the $j$-th component of the vector $\vec{\theta}_{\mathrm{opt}}$ (14).
By its very design, OS finds the numerical values of the characteristic constants $\tilde{\theta}_{1},\dots,\tilde{\theta}_{N_{\mathsf{x}}}$ minimizing the Euclidean norm $C$ (14) of the vector $\left(\log_{10}(\lambda_{1}),\log_{10}(\lambda_{2}),\dots,\log_{10}(\lambda_{N_{\mathsf{l}}})\right)$. As discussed in [6], the Euclidean norm $C$ (14) is not the unique choice for the function to minimize. By considering different objective functions, we show in [6] that the performance of OS is not affected by the particular choice of the function in use. This observation can be motivated by the equivalence of norms on the finite dimensional space $(-\infty,\infty)^{N_{\mathsf{l}}}$. Nevertheless, the availability of an explicitly computable solution makes $C$ (14) the metric of choice for Optimal Scaling.
3.2 Optimal Scaling with Constraints
Because of the similar values achieved by $\lambda_{1},\dots,\lambda_{N_{\mathsf{l}}}$, the Optimal Scaling (OS) summarized above does not allow examining regimes of $\lambda_{1},\dots,\lambda_{N_{\mathsf{l}}}$ (13) in (12) which exhibit a large difference in magnitudes. The Optimal Scaling with Constraints (OSC) proposed in [7] addresses scenarios where (at least) one coefficient is much smaller than the others, i.e., most of $\lambda_{1},\dots,\lambda_{N_{\mathsf{l}}}$ in (12) are as close as possible to $1$, while (at least) one coefficient is $\ll 1$. The method quantifies those $\theta_{1},\dots,\theta_{N_{\mathsf{x}}}\in(0,\infty)$ that lead to such a regime.
As demonstrated in [7] on the example of the Population Balance Model [8] for Latex Particles Morphology Formation [9, 10], OSC can also determine a threshold on the values of physical parameters $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$ that allow dropping the terms relative to the $\lambda_{i}\ll 1$ in the considered equations without significantly changing their solutions, but dramatically reducing computational effort.
3.2.1 Formulation
Let us denote $\mathtt{s}\in\{1,\dots,N_{\mathsf{l}}\}$ as the index of the coefficient $\lambda_{\mathtt{s}}$ being much smaller than $1$ and $U_{\mathtt{s}}\coloneqq\{1,\dots,N_{\mathsf{l}}\}\setminus\{\mathtt{s}\}$ as the set of indexes of the coefficients $\lambda_{1},\dots,\lambda_{N_{\mathsf{l}}}$ that are as close as possible to $1$. In particular, one aims to have
$\lambda_{\mathtt{s}}\ll 1,\quad\lambda_{i}\approx 1,\quad\forall i\in U_{\mathtt{s}}\coloneqq\{1,\dots,N_{\mathsf{l}}\}\setminus\{\mathtt{s}\},\quad\mbox{with}\quad\mathtt{s}\in\{1,\dots,N_{\mathsf{l}}\}.$ (16)
Given any fixed choice of $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$ and following the procedure proposed in [7], it is possible to show that the vector $\vec{\theta}=\vec{\theta}^{\star}_{\mathrm{opt}}$ targeting the regime (16) of $\lambda_{1},\dots,\lambda_{N_{\mathsf{l}}}$ (13) can be found as
$\vec{\theta}^{\star}_{\mathrm{opt}}\coloneqq\mathop{\mathrm{argmin}}_{\vec{\theta}\in(0,\infty)^{N_{\mathsf{x}}}}C_{U_{\mathtt{s}}}(\vec{\theta}),\quad C_{U_{\mathtt{s}}}(\vec{\theta})\coloneqq\sum_{i\in U_{\mathtt{s}}}\left(\log_{10}(\lambda_{i}(\tilde{\theta}_{1},\dots,\tilde{\theta}_{N_{\mathsf{x}}},\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}))-\Theta_{i}\right)^{2},$
$\mbox{where}\quad\vec{\theta}\coloneqq(\theta_{1},\dots,\theta_{N_{\mathsf{x}}})\in(0,\infty)^{N_{\mathsf{x}}},\quad\Theta_{1}=\cdots=\Theta_{N_{\mathsf{l}}}=0,$ (17)
and $C_{U_{\mathtt{s}}}$ accounts only for the coefficients $\lambda_{i}\approx 1$, as specified by $U_{\mathtt{s}}$ (16). By this means, the cost function $C_{U_{\mathtt{s}}}$ (17) measures the distance between the vector holding the orders of magnitude of coefficients $\lambda_{i}$, with $i\in U_{\mathtt{s}}$, and the corresponding vector of desired orders $\Theta_{i}=0$, with $i\in U_{\mathtt{s}}$. Assuming Property A.1 (Appendix A), provided $U=U_{\mathtt{s}}$ (16) and given any fixed $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$, the points $\vec{\theta}^{\star}_{\mathrm{opt}}$ of global minimum of $C_{U_{\mathtt{s}}}$ can be computed by means of (B.2), as follows from Theorem B.1 (Appendix B).
Then, the following features can be attained by the OSC methodology [7] and the computed $\vec{\theta}^{\star}_{\mathrm{opt}}$ (17). For any given $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$, the coefficients $\lambda_{i}$, with $i\in U_{\mathtt{s}}$ (16), result to be as close as possible to the targeted value $1$, i.e., OSC attempts to make
$\lambda_{i}(\tilde{\theta}_{1},\dots,\tilde{\theta}_{N_{\mathsf{x}}},\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}})\approx 1,\quad\mbox{when}\quad\tilde{\theta}_{j}=\vec{\theta}^{\star}_{\mathrm{opt},j}\,[\tilde{\theta}_{j}],\quad\forall i\in U_{\mathtt{s}},\quad\forall j=1,\dots,N_{\mathsf{x}},$ (18)
with $\vec{\theta}^{\star}_{\mathrm{opt},j}$ being the $j$-th component of the vector $\vec{\theta}^{\star}_{\mathrm{opt}}$ (17). Hence, the scaled model (12) accounting only for the terms relative to $\lambda_{i}$, with $i\in U_{\mathtt{s}}$ (16), inherits the features of the OS approach [6] discussed in Section 3.1.
On the other hand, the inequality
$\lambda_{\mathtt{s}}(\tilde{\theta}_{1},\dots,\tilde{\theta}_{N_{\mathsf{x}}},\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}})\ll 1,\quad\mbox{with}\quad\tilde{\theta}_{j}=\vec{\theta}^{\star}_{\mathrm{opt},j}\,[\tilde{\theta}_{j}],\quad\forall j=1,\dots,N_{\mathsf{x}},$ (19)
can be used together with (18) to identify a threshold for the values of physical parameters $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$ leading to a large difference in magnitudes between $\lambda_{\mathtt{s}}$ and all the remaining $\lambda_{i}$, with $i\in U_{\mathtt{s}}$ (16).
Finally, we remark that OS presented in Section 3.1 can be viewed as a special case of OSC discussed here. Indeed, the optimization problem (14) is equivalent to (17) with $\mathtt{s}=0$ and $U_{0}$ defined as in (16), i.e., $U_{0}\coloneqq\{1,\dots,N_{\mathsf{l}}\}\setminus\{0\}=\{1,\dots,N_{\mathsf{l}}\}$. For this reason, from now on we will focus on OSC only, with the non-negative integer $\mathtt{s}\leq N_{\mathsf{l}}$ assuming the use of OS when $\mathtt{s}=0$.
4 Generalized Optimal Scaling
Next, we intend to enrich the OSC method with an extra feature, namely the ability to reduce to its minimum the number of parameters employed in a model. With this addition, OSC can optionally control over-parametrization and should be beneficial for real-life models with a large number of physical parameters. This can be achieved by applying ideas and concepts relative to dimensional analysis and the Buckingham $\pi$ Theorem [15] in particular. From now on, we will refer to OSC with such a functionality as Generalized Optimal Scaling, or GOS.
4.1 Formulation
By requiring Properties A.1, A.2, A.3, A.4, A.5 and A.6 (Appendix A) and applying Theorems B.1, B.2, B.3, B.4, B.5, B.6 and B.7 (Appendix B), we formulate here the GOS methodology that assures all the features of OSC, while enables the minimization of the number of parameters employed.
Given Properties A.4, A.5 and A.6 of physical parameters $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$, Theorem B.3 assures that there exist only $N_{\mathsf{0}}$ independent dimensionless parameters $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}\in(0,\infty)$ being a power-law monomial of $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$. This to say that any dimensionless power-law monomial of $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$ can be written in terms of $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}$ only and, thus, $N_{\mathsf{0}}$ is the minimal number of dimensionless power-law monomials of $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$ one can consider.
By applying Properties A.1, A.2, A.3, A.5, one reveals that each coefficient $\lambda_{1},\dots,\lambda_{N_{\mathsf{l}}}$ (13) of model (12) is a dimensionless power-law monomial of $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$ and, thus, it can be written in terms of $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}$ only, as stated by (B.20) in Theorem B.4 under Properties A.1, A.2, A.3, A.4, A.5 and A.6. This implies that (9) and (12) are equivalent to a model with governing coefficients given by $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}$ only. For this reason, it is possible to reformulate the scaling approach in terms of the independent dimensionless parameters $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}$, whose number $N_{\mathsf{0}}$ is minimal.
The formulation of the GOS methodology is based on replacing $\lambda_{i}$ in (17) with the expression (B.20) provided by Theorem B.4 under Properties A.1, A.2, A.3, A.4, A.5 and A.6. Then, the optimization problem (17) is considered in terms of the vector $\vec{\alpha}\in\mathbb{R}^{N_{\mathsf{l}}\,N_{\mathsf{0}}}$ (B.39) collecting the exponents $\alpha_{i,k}$ in (B.20). As dictated by Theorem B.6 when Properties A.1, A.2, A.3, A.4, A.5 and A.6 hold, such a vector $\vec{\alpha}\in\mathbb{R}^{N_{\mathsf{l}}\,N_{\mathsf{0}}}$ (B.39) must satisfy the linear system $\textbf{S}\,\vec{\alpha}=\vec{s}$ (B.40). Then, the optimization of $\vec{\alpha}$ must take into account such a constraint, as specified in what follows.
Given any fixed choice of $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}\in(0,\infty)$, it is possible to target values of $\vec{\alpha}\in\mathbb{R}^{N_{\mathsf{l}}\,N_{\mathsf{0}}}$ (B.39) satisfying $\textbf{S}\,\vec{\alpha}=\vec{s}$ (B.40) and minimizing the deviation $C_{U_{\mathtt{s}}}:\mathbb{R}^{N_{\mathsf{l}}\,N_{\mathsf{0}}}\to[0,\infty)$ from unity of the magnitudes of $\lambda_{i}$, with $i\in U_{\mathtt{s}}\coloneqq\{1,\dots,N_{\mathsf{l}}\}\setminus\{\mathtt{s}\}$ for a fixed $\mathtt{s}\in\{0,1,\dots,N_{\mathsf{l}}\}$:
$\vec{\alpha}_{\mathrm{opt}}\coloneqq\mathop{\mathrm{argmin}}_{\begin{subarray}{c}\vec{\alpha}\in\mathbb{R}^{N_{\mathsf{l}}\,N_{\mathsf{0}}}\\ \textbf{S}\,\vec{\alpha}=\vec{s}\end{subarray}}C_{U_{\mathtt{s}}}(\vec{\alpha}),\quad C_{U_{\mathtt{s}}}(\vec{\alpha})\coloneqq\sum_{i\in U_{\mathtt{s}}}\left(\log_{10}(\lambda_{i})\right)^{2},\quad\lambda_{i}=\pi_{1}^{\alpha_{i,1}}\,\pi_{2}^{\alpha_{i,2}}\,\cdots\,\pi_{N_{\mathsf{0}}}^{\alpha_{i,N_{\mathsf{0}}}},$ (20)
where $\mathtt{s}=0$ encodes the use of Optimal Scaling (OS), while fixing $\mathtt{s}\in\{1,\dots,N_{\mathsf{l}}\}$ corresponds to the application of Optimal Scaling with Constraints (OSC).
Given $U=U_{\mathtt{s}}$, let $\boldsymbol{\Sigma}=\textbf{S}$ and $\vec{\sigma}=\vec{s}$ satisfy (B.50). Then, Theorem B.7 allows computing $\vec{\alpha}_{\mathrm{opt}}$ (20) by means of (B.49). If $\boldsymbol{\Sigma}=\textbf{S}$ and $\vec{\sigma}=\vec{s}$ do not fulfill (B.50), but such an assumption is accomplished by the row echelon form of $\textbf{S}\,\vec{\alpha}=\vec{s}$, it is possible to assign $\boldsymbol{\Sigma}$ and $\vec{\sigma}$ as the corresponding matrices of such a reduced system.
By setting $\vec{\alpha}=\vec{\alpha}_{\mathrm{opt}}$ (20), the following features can be attained by the resulting coefficients $\lambda_{1},\dots,\lambda_{N_{\mathsf{l}}}$ (13). For any fixed choice of $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}\in(0,\infty)$, the coefficients $\lambda_{i}$, with $i\in U_{\mathtt{s}}$, result to be as close as possible to the targeted value $1$, i.e., GOS attempts to make
$\lambda_{i}=\pi_{1}^{\alpha_{i,1}}\,\pi_{2}^{\alpha_{i,2}}\,\cdots\,\pi_{N_{\mathsf{0}}}^{\alpha_{i,N_{\mathsf{0}}}}\approx 1,\quad\mbox{when}\quad\vec{\alpha}\coloneqq(\alpha_{1,1},\dots,\alpha_{N_{\mathsf{l}},N_{\mathsf{0}}})^{T}=\vec{\alpha}_{\mathrm{opt}},$ (21)
for all $i\in U_{\mathtt{s}}\coloneqq\{1,\dots,N_{\mathsf{l}}\}\setminus\{\mathtt{s}\}$ and a fixed $\mathtt{s}\in\{0,1,\dots,N_{\mathsf{l}}\}$.
When $\mathtt{s}\in\{1,\dots,N_{\mathsf{l}}\}$, the inequality
$\lambda_{\mathtt{s}}=\pi_{1}^{\alpha_{\mathtt{s},1}}\,\pi_{2}^{\alpha_{\mathtt{s},2}}\,\cdots\,\pi_{N_{\mathsf{0}}}^{\alpha_{\mathtt{s},N_{\mathsf{0}}}}\ll 1,\quad\mbox{with}\quad\vec{\alpha}\coloneqq(\alpha_{1,1},\dots,\alpha_{N_{\mathsf{l}},N_{\mathsf{0}}})^{T}=\vec{\alpha}_{\mathrm{opt}},$ (22)
can be used together with (21) to identify a threshold for the values of parameters $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}$ leading to a large difference in magnitudes between $\lambda_{\mathtt{s}}$ and all the remaining $\lambda_{i}$, with $i\in U_{\mathtt{s}}$.
The objectives and features of presented scaling procedures are highlighted in Table 1.
Optimal Scaling with Constraints Generalized Optimal Scaling
Objectives Finds coefficients $\lambda_{1},\dots,\lambda_{N_{\mathsf{l}}}$ in (12) that meet conditions: $\lambda_{\mathtt{s}}\ll 1$, $\lambda_{i}\approx 1$, $\forall i\neq\mathtt{s}\in\{1,\dots,N_{\mathsf{l}}\}$ Finds coefficients $\lambda_{1},\dots,\lambda_{N_{\mathsf{l}}}$ in (12) that meet conditions: $\lambda_{\mathtt{s}}\ll 1$, $\lambda_{i}\approx 1$, $\forall i\neq\mathtt{s}\in\{1,\dots,N_{\mathsf{l}}\}$
Controls over-parametrization
Employed Assumptions Property A.1 Properties A.1, A.2, A.3, A.4, A.5 and A.6
Features Provides the conditions (18)-(19) that allow discarding the terms in (12) related to $\lambda_{\mathtt{s}}\ll 1$, without significantly modifying the solutions to (12)
The model (12) with $\lambda_{\mathtt{s}}=0$ shows minimal variability $\mathop{\mathrm{max}}_{i\neq\mathtt{s}}\lambda_{i}/\mathop{\mathrm{min}}_{i\neq\mathtt{s}}\lambda_{i}$ and reduced round-off errors
Inherits all the features of Optimal Scaling with Constraints, with the conditions (18)-(19) replaced by (21)-(22)
Employs the minimal number of governing parameters
Procedure Given any fixed choice of physical parameters $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$, finds $\vec{\theta}\coloneqq(\theta_{1},\dots,\theta_{N_{\mathsf{x}}})\in(0,\infty)^{N_{\mathsf{x}}}$ that minimizes the deviation $C_{U_{\mathtt{s}}}$ (17) from unity of the magnitudes of $\lambda_{i}$, $\forall i\in U_{\mathtt{s}}\coloneqq\{1,\dots,N_{\mathsf{l}}\}\setminus\{\mathtt{s}\}$ Given any fixed choice of $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}\in(0,\infty)$, finds $\vec{\alpha}$ (B.39) that satisfies (B.40) and minimizes the deviation $C_{U_{\mathtt{s}}}$ (20) from unity of the magnitudes of $\lambda_{i}$, $\forall i\in U_{\mathtt{s}}\coloneqq\{1,\dots,N_{\mathsf{l}}\}\setminus\{\mathtt{s}\}$
Realization The points $\vec{\theta}$ of global minimum of $C_{U_{\mathtt{s}}}$ (17) can be computed by means of (B.2) provided $U=U_{\mathtt{s}}$, as follows from Theorem B.1 With $U=U_{\mathtt{s}}$, $\boldsymbol{\Sigma}=\textbf{S}$, $\vec{\sigma}=\vec{s}$ (B.40) and assuming (B.50), Theorem B.7 allows computing $\vec{\alpha}=\vec{\alpha}_{\mathrm{opt}}$ (20) by means of (B.49)
Plugs $\vec{\alpha}=\vec{\alpha}_{\mathrm{opt}}$ (20) into (B.38), then solves for $\vec{t}$ (B.24) the system (B.25) and uses the found $\vec{t}$ to compute $\tilde{\theta}_{1},\dots,\tilde{\theta}_{N_{\mathsf{x}}}$ as in (A.3)
Table 1: Summary and features of Optimal Scaling with Constraints [7] and Generalized Optimal Scaling (GOS). The Optimal Scaling [6] is equivalent to GOS with $\mathtt{s}=0$.
5 A Case Study: the classical projectile model
Aiming to test the scaling techniques discussed above, we apply them to the projectile model introduced in Section 2.2.
First, we use the Theorems from Appendix B to formulate the projectile model with the minimal number of parameters. Properties A.4, A.5 and A.6 are satisfied by the $N_{\mathsf{p}}=4$ physical parameters $\tilde{p}_{1},\tilde{p}_{2},\tilde{p}_{3},\tilde{p}_{4}$ (6), with
$[\tilde{p}_{1}]=[\tilde{g}]=\mathrm{m}\,\mathrm{s}^{-2},\quad[\tilde{p}_{2}]=[\tilde{R}]=\mathrm{m},\quad[\tilde{p}_{3}]=[\tilde{h}_{0}]=\mathrm{m},\quad[\tilde{p}_{4}]=[\tilde{v}_{0}]=\mathrm{m}\,\mathrm{s}^{-1},$ (23)
being $\mathrm{m}$ meters and $\mathrm{s}$ seconds. Then, Theorem B.3 guarantees that there exist only $N_{\mathsf{0}}=N_{\mathsf{p}}-\mathrm{rank}(\textbf{M})=2$ independent dimensionless parameters, where M is the matrix
$\textbf{M}=\begin{pmatrix}1&1&1&1\\ -2&0&0&-1\end{pmatrix},$ (24)
and the $N_{\mathsf{0}}=2$ independent dimensionless parameters can be computed as in (B.12), i.e.,
$\pi_{1}=\tilde{g}^{-1/2}\,\tilde{R}^{-1/2}\,\tilde{v}_{0},\quad\pi_{2}=\tilde{R}^{-1}\,\tilde{h}_{0},\quad\pi_{1},\pi_{2}\in(0,\infty).$ (25)
Properties A.1 and A.2 are fulfilled by $N_{\mathsf{l}}=4$ coefficients $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}$ (5), with
$\tilde{\kappa}_{1}=\tilde{p}_{1}=\tilde{g},\quad\tilde{\kappa}_{2}=\tilde{p}_{2}^{-1}=\tilde{R}^{-1},\quad\tilde{\kappa}_{3}=\tilde{p}_{3}=\tilde{h}_{0},\quad\tilde{\kappa}_{4}=\tilde{p}_{4}=\tilde{v}_{0},$ (26)
and
$c_{1,1}=2,$ $\quad c_{1,2}=-1,$
$c_{2,1}=0,$ $\quad c_{2,2}=1,$
$c_{3,1}=0,$ $\quad c_{3,2}=-1,$
$c_{4,1}=1,$ $\quad c_{4,2}=-1.$ (27)
Then, by also assuming Property A.3 for $\tilde{\theta}_{1},\tilde{\theta}_{2}$ (3), i.e.,
$\tilde{\theta}_{1}$ $=\tilde{p}_{1}^{t_{1,1}}\,\tilde{p}_{2}^{t_{2,1}}\,\tilde{p}_{3}^{t_{3,1}}\,\tilde{p}_{4}^{t_{4,1}}=\tilde{g}^{t_{1,1}}\,\tilde{R}^{t_{2,1}}\,\tilde{h}_{0}^{t_{3,1}}\,\tilde{v}_{0}^{t_{4,1}},$
$\tilde{\theta}_{2}$ $=\tilde{p}_{1}^{t_{1,2}}\,\tilde{p}_{2}^{t_{2,2}}\,\tilde{p}_{3}^{t_{3,2}}\,\tilde{p}_{4}^{t_{4,2}}=\tilde{g}^{t_{1,2}}\,\tilde{R}^{t_{2,2}}\,\tilde{h}_{0}^{t_{3,2}}\,\tilde{v}_{0}^{t_{4,2}},$ (28)
Theorem B.4 guarantees that the coefficients $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}$ (5) can be written as power-law monomials of $\pi_{1},\pi_{2}$ (25) only, i.e.,
$\lambda_{1}=\pi_{1}^{\alpha_{1,1}}\,\pi_{2}^{\alpha_{1,2}},\quad\lambda_{2}=\pi_{1}^{\alpha_{2,1}}\,\pi_{2}^{\alpha_{2,2}},\quad\lambda_{3}=\pi_{1}^{\alpha_{3,1}}\,\pi_{2}^{\alpha_{3,2}},\quad\lambda_{4}=\pi_{1}^{\alpha_{4,1}}\,\pi_{2}^{\alpha_{4,2}}.$ (29)
Finally, Theorem B.6 indicates under Properties A.1, A.2, A.3, A.4, A.5 and A.6 that the vector
$\vec{\alpha}=\begin{pmatrix}\alpha_{1,1}&\alpha_{1,2}&\alpha_{2,1}&\alpha_{2,2}&\alpha_{3,1}&\alpha_{3,2}&\alpha_{4,1}&\alpha_{4,2}\end{pmatrix}^{T}\in\mathbb{R}^{8}$ (30)
of exponents in (29) must satisfy the linear system $\textbf{S}\,\vec{\alpha}=\vec{s}$ (B.40), which, for the projectile model, reads in row echelon form as
$\begin{pmatrix}-1/2&0&1/2&0&0&0&1&0\\ 0&-1/2&0&1/2&0&0&0&1\\ 0&0&1&0&1&0&0&0\\ 0&0&0&1&0&1&0&0\end{pmatrix}\vec{\alpha}=\begin{pmatrix}1\\ 0\\ 0\\ 1\end{pmatrix}.$ (31)
5.1 Optimally Scaled Coefficients
We compute the values of coefficients $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}$ (5) using the Generalized Optimal Scaling approach presented in Section 4. In particular, we provide and discuss the values of coefficients $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}$ (29) when the vector $\vec{\alpha}$ (30) of exponents in (29) is optimized according to (20). In this case, $U=U_{\mathtt{s}}\coloneqq\{1,2,3,4\}\setminus\{\mathtt{s}\}$, where $\mathtt{s}=0$ means the use of OS, while fixing $\mathtt{s}\in\{1,2,3,4\}$ corresponds to the application of OSC.
Theorem B.7 allows solving the optimization problem (20) for any fixed choice of $\pi_{1},\pi_{2}$ (25), since one can set (31) as the constraint $\boldsymbol{\Sigma}\,\vec{\alpha}=\vec{\sigma}$ (B.48) to satisfy the assumption (B.50). In particular, the system (B.49) can be solved to find optimal values of $\vec{\alpha}$ (30) for plugging into (29). Then, the system (B.49) reads as
$\begin{pmatrix}\boldsymbol{\Sigma}&\textbf{0}_{4\times 4}\\ \textbf{G}&\boldsymbol{\Sigma}^{T}\end{pmatrix}\begin{pmatrix}\vec{\alpha}\\ \vec{\nu}\end{pmatrix}=\begin{pmatrix}\vec{\sigma}\\ \textbf{0}_{8\times 1}\end{pmatrix},$ (32)
with $\vec{\alpha}\in\mathbb{R}^{8}$ given by (30), $\vec{\nu}=(\nu_{1},\nu_{2},\nu_{3},\nu_{4})^{T}\in\mathbb{R}^{4}$,
$\boldsymbol{\Sigma}=\begin{pmatrix}-1/2&0&1/2&0&0&0&1&0\\ 0&-1/2&0&1/2&0&0&0&1\\ 0&0&1&0&1&0&0&0\\ 0&0&0&1&0&1&0&0\end{pmatrix},\quad\vec{\sigma}=\begin{pmatrix}1\\ 0\\ 0\\ 1\end{pmatrix},$ (33)
$\textbf{G}=\begin{pmatrix}\chi_{\mathtt{s}}(1)\,\textbf{P}&\textbf{0}_{2\times 2}&\textbf{0}_{2\times 2}&\textbf{0}_{2\times 2}\\ \textbf{0}_{2\times 2}&\chi_{\mathtt{s}}(2)\,\textbf{P}&\textbf{0}_{2\times 2}&\textbf{0}_{2\times 2}\\ \textbf{0}_{2\times 2}&\textbf{0}_{2\times 2}&\chi_{\mathtt{s}}(3)\,\textbf{P}&\textbf{0}_{2\times 2}\\ \textbf{0}_{2\times 2}&\textbf{0}_{2\times 2}&\textbf{0}_{2\times 2}&\chi_{\mathtt{s}}(4)\,\textbf{P}\end{pmatrix},\quad\textbf{P}=\begin{pmatrix}2(\log_{10}(\pi_{1}))^{2}&2\log_{10}(\pi_{1})\log_{10}(\pi_{2})\\ 2\log_{10}(\pi_{1})\log_{10}(\pi_{2})&2(\log_{10}(\pi_{2}))^{2}\end{pmatrix},$ (34)
where $\chi_{\mathtt{s}}(i)=1$ if $i\neq\mathtt{s}$, $\chi_{\mathtt{s}}(i)=0$ if $i=\mathtt{s}$, and $\textbf{0}_{m\times n}$ is a matrix with $m$ rows and $n$ columns containing all zeros. By plugging the solutions to (32)-(34) into (29), one obtains
$\lambda_{1}=\pi_{1}^{-4/11}\,\pi_{2}^{1/11},$ $\lambda_{2}=\pi_{1}^{2/11}\,\pi_{2}^{5/11},$ $\lambda_{3}=\pi_{1}^{-2/11}\,\pi_{2}^{6/11},$ $\lambda_{4}=\pi_{1}^{8/11}\,\pi_{2}^{-2/11},$ $\mbox{if}\quad\mathtt{s}=0,$ (35)
$\lambda_{1}=\pi_{1}^{-2}\,\pi_{2}^{1/2},$ $\lambda_{2}=\pi_{2}^{1/2},$ $\lambda_{3}=\pi_{2}^{1/2},$ $\lambda_{4}=1,$ $\mbox{if}\quad\mathtt{s}=1,$ (36)
$\lambda_{1}=\pi_{1}^{-1/3}\,\pi_{2}^{1/6},$ $\lambda_{2}=\pi_{1}^{1/3}\,\pi_{2}^{5/6},$ $\lambda_{3}=\pi_{1}^{-1/3}\,\pi_{2}^{1/6},$ $\lambda_{4}=\pi_{1}^{2/3}\,\pi_{2}^{-1/3},$ $\mbox{if}\quad\mathtt{s}=2,$ (37)
$\lambda_{1}=\pi_{1}^{-1/3},$ $\lambda_{2}=\pi_{1}^{1/3},$ $\lambda_{3}=\pi_{1}^{-1/3}\,\pi_{2},$ $\lambda_{4}=\pi_{1}^{2/3},$ $\mbox{if}\quad\mathtt{s}=3,$ (38)
$\lambda_{1}=1,$ $\lambda_{2}=\pi_{2}^{1/2},$ $\lambda_{3}=\pi_{2}^{1/2},$ $\lambda_{4}=\pi_{1}\,\pi_{2}^{-1/4},$ $\mbox{if}\quad\mathtt{s}=4.$ (39)
The coefficients $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}$ (35)-(39) are depicted in Figure 1 as functions of the parameters $\pi_{1},\pi_{2}$ (25). For all tested $\mathtt{s}=0,1,2,3,4$, GOS identifies the values of $\pi_{1},\pi_{2}$ which satisfy the regime (21)-(22). Such values lie in the areas indicated in Figure 1 by magenta dashed lines.
(a) Optimally scaled coefficients $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}$ (35) with $\mathtt{s}=0$.
(b) Optimally scaled coefficients $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}$ (36) with $\mathtt{s}=1$.
(c) Optimally scaled coefficients $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}$ (37) with $\mathtt{s}=2$.
(d) Optimally scaled coefficients $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}$ (38) with $\mathtt{s}=3$.
(e) Optimally scaled coefficients $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}$ (39) with $\mathtt{s}=4$.
Figure 1: Coefficients (35)-(39) provided by GOS as functions of $\pi_{1},\pi_{2}$ (25). The regime (21)-(22) is achieved in the areas delimited by magenta dashed lines.
5.2 Generalized Optimal Scaling $(\mathtt{s}=0)$: Results
Next, we illustrate the advantages provided by the coefficients $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}$ (35) resulting from application of Generalized Optimal Scaling with $\mathtt{s}=0$ to the projectile model.
Table 1 states that such coefficients allow minimizing the variability $\mathop{\mathrm{max}}_{i}\lambda_{i}/\mathop{\mathrm{min}}_{i}\lambda_{i}$. In order to verify such a claim, we compare the values achieved by coefficients (35) with the obtained ones by the following approach suggested in the literature [13, 14]. Provided the values of physical parameters $\tilde{p}_{1},\tilde{p}_{2},\tilde{p}_{3},\tilde{p}_{4}$ (6), it is possible to impose $N_{\mathsf{x}}=2$ out of the $N_{\mathsf{l}}=4$ coefficients $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}$ (5) to be equal to $1$. Then, one aims to solve the resulting system of equations with respect to $\tilde{\theta}_{1}$ and $\tilde{\theta}_{2}$. Finally, the corresponding values of $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}$ (5) are obtained by plugging the computed $\tilde{\theta}_{1}$ and $\tilde{\theta}_{2}$ into (5).
Aiming to perform such a comparison, we set $\tilde{g}=9.81\,\mathrm{m}\,\mathrm{s}^{-2}$, $\tilde{R}=6.3781\times 10^{6}\,\mathrm{m}$, $\tilde{h}_{0}=1\,\mathrm{m}$ and $\tilde{v}_{0}=1\,\mathrm{m}\,\mathrm{s}^{-1}$, being $\tilde{g}$ the gravitational acceleration on the planet Earth and $\tilde{R}$ the radius of the Earth. Though the chosen values of initial height $\tilde{h}_{0}$ and initial speed $\tilde{v}_{0}$ have no particular physical meaning, they are reasonable for a ball thrown vertically upward from the surface of the Earth. The comparison of $\lambda_{i}$ values for different scaling methods and the corresponding variability $\mathop{\mathrm{max}}_{i}\lambda_{i}/\mathop{\mathrm{min}}_{i}\lambda_{i}$ are shown in 2(a) and 2(b), respectively. In the performed experiment, GOS allows reducing the variability of coefficients $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}$ by at least two more orders of magnitude compared to any of other traditional scaling approaches, as illustrated by 2(b).
(a) Coefficients $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}$ (5) are shown by colored symbols, while coefficients $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}$ (35) are depicted with black circles.
(b) Variability $\mathop{\mathrm{max}}_{i}\lambda_{i}/\mathop{\mathrm{min}}_{i}\lambda_{i}$, with $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}$ (5) and (35) shown by colored symbols and black circles, respectively.
Figure 2: Values attained by coefficients $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}$ (5) and (35) (2(a)) and corresponding variability $\mathop{\mathrm{max}}_{i}\lambda_{i}/\mathop{\mathrm{min}}_{i}\lambda_{i}$ (2(b)), when $\tilde{g}=9.81\,\mathrm{m}\,\mathrm{s}^{-2}$, $\tilde{R}=6.3781\times 10^{6}\,\mathrm{m}$, $\tilde{h}_{0}=1\,\mathrm{m}$, $\tilde{v}_{0}=1\,\mathrm{m}\,\mathrm{s}^{-1}$ and $\pi_{1},\pi_{2}$ (25). In both 2(a) and 2(b), the black circles provide the numerical values obtained by Generalized Optimal Scaling (GOS) with $\mathtt{s}=0$, i.e., the coefficients $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}$ (35). 2(a) and 2(b) also show the values achieved by $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}$ (5) if $N_{\mathsf{x}}=2$ out of the $N_{\mathsf{l}}=4$ are imposed to be equal to $1$, as indicated by the legend in 2(a) and 2(b). The Figuresdo not show the data corresponding to $\lambda_{2}=\lambda_{3}=1$, since it leads to $\tilde{p}_{2}^{-1}\,\tilde{\theta}_{2}=\tilde{p}_{3}\,\tilde{\theta}_{2}^{-1}=1$ and there are no $\tilde{\theta}_{1}$ and $\tilde{\theta}_{2}$ solving such a system of equations.
The minimal variability of the coefficients $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}$ is beneficial for reducing round-off errors, as claimed in Table 1. In order to illustrate the goodness of the GOS approach for numerical computations, we perform the following experiment.
The Matlab solver $\mathrm{\it{ode45}}$ [16] is employed to compute the numerical solution $\varphi=\varphi_{n}$ to the Ordinary Differential Equation (4) in the time interval $[0,T/\theta_{1}]$. The subscript $n$ indicates that the numerical values of coefficients $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}$ and any computed quantity at each time step are rounded to $n$ decimal places. The relative error is estimated as
[…]
where $\varphi_{\mathrm{ref}}$ is the numerical solution to the Ordinary Differential Equation (4) with $\lambda_{1}=g$, $\lambda_{2}=R^{-1}$, $\lambda_{3}=h_{0}$, $\lambda_{4}=v_{0}$, i.e., $\theta_{1}=\theta_{2}=1$, that has been taken as a reference in the time interval $[0,T]$. Such a solution $\varphi_{\mathrm{ref}}$ is obtained by using the Matlab solver $\mathrm{\it{ode45}}$ [16] with the default 16 significant digits of precision for numerical computations. A relative error tolerance of $10^{-5}$ and an absolute error tolerance of $10^{-8}$ have been required for all the computations performed by $\mathrm{\it{ode45}}$ [16], being $2$ orders of magnitude smaller than the default tolerances. The Simpson’s quadrature rule [17] has been used to compute the integral in (40).
3(a) shows the reference solution $\varphi_{\mathrm{ref}}$, while 3(b) compares the error $e_{n}$ (40) when $\varphi_{n}$ is computed with several choices of coefficients $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}$, as indicated in the legend of 3(b). The colored symbols denote the error corresponding to $\varphi_{n}$ calculated by using the numerical values of $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}$ obtained by imposing $N_{\mathsf{x}}=2$ out of the $N_{\mathsf{l}}=4$ coefficients $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}$ (5) to be equal to $1$, as detailed above. It is possible to notice that the numerical values of $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}$ (35) provided by the GOS approach allow reducing the error $e_{n}$ (black circles) up to two more orders of magnitude, compared to the other traditional scaling approaches.
(a) Reference solution $\varphi=\varphi_{\mathrm{ref}}$ to (4).
(b) Error $e_{n}$ (40).
Figure 3: 3(a) shows the numerical solution $\varphi=\varphi_{\mathrm{ref}}$ to the Ordinary Differential Equation (4) in the time interval $[0,T]$, with $\lambda_{1}=g$, $\lambda_{2}=R^{-1}$, $\lambda_{3}=h_{0}$, $\lambda_{4}=v_{0}$ and $T=700$, being $\tilde{g}=9.81\,\mathrm{m}\,\mathrm{s}^{-2}$, $\tilde{R}=6.3781\times 10^{6}\,\mathrm{m}$, $\tilde{h}_{0}=1\,\mathrm{m}$, $\tilde{v}_{0}=1\,\mathrm{m}\,\mathrm{s}^{-1}$. Given the shown $\varphi_{\mathrm{ref}}$ and $\tilde{g}$, $\tilde{R}$, $\tilde{h}_{0}$, $\tilde{v}_{0}$ as above, 3(b) depicts the error $e_{n}$ (40) corresponding to several choices of the coefficients $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}$, as indicated in the legend. The black circles show the error made by using $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}$ (35) with $\pi_{1},\pi_{2}$ (25), as provided by Generalized Optimal Scaling (GOS) with $\mathtt{s}=0$. On the other hand, the colored symbols indicate the error corresponding to numerical values of $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}$ obtained by imposing $N_{\mathsf{x}}=2$ out of the $N_{\mathsf{l}}=4$ coefficients $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}$ (5) to be equal to $1$. 3(b) does not show the data corresponding to $\lambda_{2}=\lambda_{3}=1$, since it leads to $\tilde{p}_{2}^{-1}\,\tilde{\theta}_{2}=\tilde{p}_{3}\,\tilde{\theta}_{2}^{-1}=1$ and there are no $\tilde{\theta}_{1}$ and $\tilde{\theta}_{2}$ solving such a system of equations.
5.3 Generalized Optimal Scaling $(\mathtt{s}=1,2,3,4)$: Results
Finally, we show that, in the case of the projectile model, the Generalized Optimal Scaling method is able to provide a threshold on values of parameters that allow discarding terms in the equation (4) without significantly modifying its solution.
By applying the GOS approach with a fixed $\mathtt{s}\in\{1,2,3,4\}$ to the projectile model, one aims to identify a quantitative criterion for allowing the approximation of $\varphi$ in (4) by the solutions $\varphi=\varphi_{\mathtt{s}}$ of (4) with $\lambda_{\mathtt{s}}=0$, i.e.,
$\varphi_{\mathtt{s}}(\tau)=\begin{cases}\lambda_{4}\,\tau\,+\,\lambda_{3},\quad\forall\tau\geq 0,&\mbox{if}\quad\mathtt{s}=1,\\ (-\lambda_{1}/2)\,\tau^{2}\,+\,\lambda_{4}\,\tau\,+\,\lambda_{3},\quad\forall\tau\geq 0,&\mbox{if}\quad\mathtt{s}=2,\\ \mbox{solution to}\quad\frac{d^{2}\varphi_{\mathtt{s}}}{d\tau^{2}}=-\frac{\lambda_{1}}{(1+\lambda_{2}\,\varphi_{\mathtt{s}})^{2}},\quad\forall\tau\geq 0,\quad\varphi_{\mathtt{s}}(0)=0,\quad\frac{d\varphi_{\mathtt{s}}}{d\tau}(0)=\lambda_{4},&\mbox{if}\quad\mathtt{s}=3,\\ \mbox{solution to}\quad\frac{d^{2}\varphi_{\mathtt{s}}}{d\tau^{2}}=-\frac{\lambda_{1}}{(1+\lambda_{2}\,\varphi_{\mathtt{s}})^{2}},\quad\forall\tau\geq 0,\quad\varphi_{\mathtt{s}}(0)=\lambda_{3},\quad\frac{d\varphi_{\mathtt{s}}}{d\tau}(0)=0,&\mbox{if}\quad\mathtt{s}=4.\end{cases}$ (41)
In order to quantify the goodness of approximation, we first set the coefficients $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}$ as in (36), (37), (38), (39) for $\mathtt{s}=1,2,3,4$, respectively, and then, estimate the approximation error as
[…]
$T_{0}$$\coloneqq\begin{cases}\mathop{\mathrm{min}}\left\{\tau\geq 0\,:\,\left|1\,+\,\lambda_{2}\,\varphi_{\mathrm{ref}}(\tau)\right|\leq 10^{-2}\right\},&\mbox{if}\quad\mathtt{s}=1,2,\\ \mathop{\mathrm{min}}\left\{\tau\geq 0\,:\,\left|1\,+\,\lambda_{2}\,\varphi_{\mathtt{s}}(\tau)\right|\leq 10^{-2}\,\lor\,\left|1\,+\,\lambda_{2}\,\varphi_{\mathrm{ref}}(\tau)\right|\leq 10^{-2}\right\},&\mbox{if}\quad\mathtt{s}=3,4,\\ \end{cases}$ (42)
where $\varphi_{\mathtt{s}}(\tau)$ is the approximating solution (41), while $\varphi_{\mathrm{ref}}(\tau)$ is the numerical solution to (4) provided by the Matlab solver $\mathrm{\it{ode45}}$ [16], that has been taken as a reference since it is the solution one aims to approximate. If $\mathtt{s}=3,4$, the approximating solution (41) is also numerically computed by means of the Matlab solver $\mathrm{\it{ode45}}$ [16]. A relative error tolerance of $10^{-6}$ and an absolute error tolerance of $10^{-9}$ have been required for all the computations performed by $\mathrm{\it{ode45}}$ [16], i.e., $3$ orders of magnitude smaller than the default tolerances. The Simpson’s quadrature rule [17] has been used to compute the integrals in (42). The time $T$ in (42) is chosen with the aim of stopping the numerical integration just before singularities are attained in (4) or (41). Such irregularities appear when the computed solutions of (4) or (41) are negative, i.e., equal to $-1/\lambda_{2}<0$. In other words, the time interval $[0,T]$ in (42) allows investigating the dynamics of the projectile further than it reaches the ground level. The resulting error $\epsilon_{\mathtt{s}}(T)$ (42) is shown in Figure 4 as a function of $\pi_{1}$ and $\pi_{2}$ (25).
As explained in Section 4.1, GOS identifies (21)-(22) as conditions for a large difference in magnitudes between $\lambda_{\mathtt{s}}$ and all the remaining $\lambda_{i}$, with $i\neq\mathtt{s}\in\{1,\dots,N_{\mathsf{l}}\}$. Such requirements can be written for some $\delta\in(-\infty,-1]$ and $\gamma\in(0,1]$ as
$\lambda_{\mathtt{s}}\leq 10^{\delta}\quad\wedge\quad 10^{-\gamma}<\lambda_{i}<10^{\gamma},\quad\forall i\neq\mathtt{s}.$ (43)
By using (36), (37), (38), (39) for $\mathtt{s}=1,2,3,4$, respectively, the inequalities (43) give
$\lambda_{\mathtt{s}}=\pi_{1}^{-2}\,\pi_{2}^{1/2}\leq 10^{\delta}$$\wedge$ $-2\gamma<\log_{10}(\pi_{2})<2\gamma,$ $\mbox{if}\quad\mathtt{s}=1,$
$\lambda_{\mathtt{s}}=\pi_{1}^{1/3}\,\pi_{2}^{5/6}\leq 10^{\delta}$$\wedge$ $-3\gamma<\log_{10}(\pi_{2})-2\log_{10}(\pi_{1})<3\gamma,$ $\mbox{if}\quad\mathtt{s}=2,$
$\lambda_{\mathtt{s}}=\pi_{1}^{-1/3}\,\pi_{2}\leq 10^{\delta}$$\wedge$ $-3\gamma<2\log_{10}(\pi_{1})<3\gamma,$ $\mbox{if}\quad\mathtt{s}=3,$
$\lambda_{\mathtt{s}}=\pi_{1}\,\pi_{2}^{-1/4}\leq 10^{\delta}$$\wedge$ $-2\gamma<\log_{10}(\pi_{2})<2\gamma,$ $\mbox{if}\quad\mathtt{s}=4,$ (44)
with $\delta\in(-\infty,-1]$ and $\gamma\in(0,1]$. Figure 4 shows that requiring (44) with $\delta=-2$ and $\gamma=1/2$ guarantees a relative error of at most $10\%$, i.e., $\epsilon_{\mathtt{s}}(T)\leq 10^{-1}$, for all simulated intervals of time $T$ and all $\mathtt{s}=1,2,3,4$. In Figure 4, the threshold $\epsilon_{\mathtt{s}}(T)=10^{-1}$ is highlighted by white solid lines, while the conditions (44) with $\delta=-2$ and $\gamma=1/2$ are satisfied by the values of $\pi_{1},\pi_{2}$ lying in the areas delimited by magenta dashed lines. Finally, as clearly demonstrated by Figure 4, the approximation by an asymptotic model is improving, i.e., the value of $\epsilon_{\mathtt{s}}(T)$ is decreasing for all tested $\mathtt{s}$ and $T$, by choosing a smaller $\delta\in(-\infty,-2]$ at the fixed $\gamma=1/2$ in (44) to obtain the ranges of appropriate physical parameters for such a model.
6 Conclusions & Discussion
We reviewed methodologies for scaling of dimensional models and presented a novel approach which we called Generalized Optimal Scaling, or GOS. GOS inherits all the features of its predecessors, i.e., Optimal Scaling (OS) and Optimal Scaling with Constraints (OSC), including the ability to quantify conditions for asymptotic models and, in addition, provides control of over-parametrization - the property beneficial for complex multiscale models. Both OS and OSC can be recovered from GOS with the special choices of the GOS’ parameter $\mathtt{s}$. Such a non-negative integer $\mathtt{s}$ is the index of the coefficient $\lambda_{\mathtt{s}}$ vanishing to 00 when an asymptotic model is considered. The GOS method was rigorously formulated and carefully tested on the classical projectile model.
By setting $\mathtt{s}=0$ in GOS (see Section 4), one makes use of the Optimal Scaling methodology [6] presented in Section 3.1. For the projectile model, we demonstrated that OS allows reducing the variability of the involved coefficients by at least two more orders of magnitude compared to other scaling approaches traditionally applied to this model. Such a reduction is beneficial for diminishing round-off and numerical errors. Indeed, we showed that the coefficients provided by the Optimal Scaling approach allow reducing round-off errors by two more orders of magnitude compared to the traditional scaling techniques.
When GOS is applied with $\mathtt{s}\geq 1$ (see Section 4), one performs the Optimal Scaling with Constraints [7] explained in Section 3.2. Such an approach provides a threshold on values of parameters that allow discarding terms in the considered equations, without significantly modifying the solutions. Thus, one can drop some terms and consider a simpler model, if the provided criterion is met. For the projectile model, OSC is able to identify values of parameters yielding to a relative approximation error of at most $10\%$ for all simulated intervals of time and all $\mathtt{s}\geq 1$.
A natural future step is an incorporation of GOS in an automated tool to streamline model reduction and scaling, optimizing parameter selection before further modeling steps. This would facilitate the use of GOS in complex, high-dimensional models, even for users with less mathematical expertise.
Figure 4: Error $\epsilon_{\mathtt{s}}(T)$ (42) and coefficient $\lambda_{\mathtt{s}}$ (44) as functions of $\pi_{1},\pi_{2}$ (25). Along the white solid and dotted lines, the error $\epsilon_{\mathtt{s}}(T)$ is equal to $10^{-1}$ and $10^{-2}$, respectively. The inequalities (44) are met with $\delta=-2$ and $\gamma=1/2$ in the areas delimited by magenta dashed lines.
Acknowledgements
This research was supported by MICIU/AEI/10.13039/501100011033 and by ERDF A way for Europe under Grant PID2022-136585NB-C22; funded by MICIU/AEI/10.13039/501100011033 and cofunded by the European Union under Grant PID2023-146764NB-I00; financed by the Basque Government through ELKARTEK Programme under Grant KK-2024/00062 and through the BERC 2022-2025 program.
[…]
The authors acknowledge the financial support received from the grant BCAM-IKUR, funded by the Basque Government, by the IKUR Strategy and by the European Union NextGenerationEU/PRTR.
This work has been possible thanks to the support of the computing infrastructure of Barcelona Supercomputing Center (RES, QHS-2025-1-0027) and DIPC Computer Center.
Appendix A Properties
The Optimal Scaling approaches presented in this paper rely on the Properties listed below. Such assumptions give rise to the Theorems formulated in Appendix B to constitute the tools for Optimal Scaling procedures.
Property A.1.
Each of the dimensionless coefficients $\lambda_{1},\dots,\lambda_{N_{\mathsf{l}}}$ (13) can be written as a power-law monomial of $\tilde{\theta}_{1},\dots,\tilde{\theta}_{N_{\mathsf{x}}}$ multiplied by a factor $\tilde{\kappa}_{i}$:
$\lambda_{i}=\tilde{\kappa}_{i}\,\tilde{\theta}_{1}^{c_{i,1}}\,\tilde{\theta}_{2}^{c_{i,2}}\,\cdots\,\tilde{\theta}_{N_{\mathsf{x}}}^{c_{i,N_{\mathsf{x}}}},\quad\tilde{\kappa}_{i}=\tilde{\kappa}_{i}(\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}),\quad p_{1},\dots,p_{N_{\mathsf{p}}}\in(0,\infty),$
$\kappa_{i},\theta_{j}\in(0,\infty),\quad c_{i,j}\in\mathbb{R},\quad\forall i=1,\dots,N_{\mathsf{l}},\quad\forall j=1,\dots,N_{\mathsf{x}},$ (A.1)
where each $\tilde{\kappa}_{i}$ is an explicit function of $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$ only, and all the exponents $c_{i,j}$ in (A.1) are explicitly known real numbers.
Property A.2.
Each of the factors $\tilde{\kappa}_{1},\dots,\tilde{\kappa}_{N_{\mathsf{l}}}$ in (A.1) can be written as a power-law monomial of the physical parameters $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$ only:
$\tilde{\kappa}_{i}=\tilde{p}_{1}^{d_{i,1}}\,\tilde{p}_{2}^{d_{i,2}}\,\cdots\,\tilde{p}_{N_{\mathsf{p}}}^{d_{i,N_{\mathsf{p}}}},\quad d_{i,j}\in\mathbb{R},\quad\forall i=1,\dots,N_{\mathsf{l}},\quad\forall j=1,\dots,N_{\mathsf{p}},$ (A.2)
with all the exponents $d_{i,j}$ in (A.2) to be explicitly known real numbers.
Property A.3.
Each of the characteristic constants $\tilde{\theta}_{1},\dots,\tilde{\theta}_{N_{\mathsf{x}}}$ can be written as a power-law monomial of the physical parameters $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$ only:
$\tilde{\theta}_{j}=\tilde{p}_{1}^{t_{1,j}}\,\tilde{p}_{2}^{t_{2,j}}\,\cdots\,\tilde{p}_{N_{\mathsf{p}}}^{t_{N_{\mathsf{p}},j}},\quad t_{i,j}\in\mathbb{R},\quad\forall i=1,\dots,N_{\mathsf{p}},\quad\forall j=1,\dots,N_{\mathsf{x}},$ (A.3)
where the exponents $t_{i,j}$ in (A.3) are unknowns of scaling approaches.
Property A.4.
The physical parameters $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$ are independent, i.e., $\tilde{p}_{1}^{\beta_{1}}\,\tilde{p}_{2}^{\beta_{2}}\,\cdots\,\tilde{p}_{N_{\mathsf{p}}}^{\beta_{N_{\mathsf{p}}}}=1\Leftrightarrow\beta_{1}=\beta_{2}=\cdots=\beta_{N_{\mathsf{p}}}=0$.
Property A.5.
The physical parameters $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$ assume strictly positive numerical values $p_{1},\dots,p_{N_{\mathsf{p}}}\in(0,\infty)$.
Property A.6.
The dimension of each $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$ is not equal to $1$ and given by a power-law monomial of $N_{\mathsf{u}}$ independent units $\mathrm{u}_{1},\dots,\mathrm{u}_{N_{\mathsf{u}}}$, i.e.,
$[\tilde{p}_{j}]=\mathrm{u}_{1}^{m_{1,j}}\,\mathrm{u}_{2}^{m_{2,j}}\,\cdots\,\mathrm{u}_{N_{\mathsf{u}}}^{m_{N_{\mathsf{u}},j}}\neq 1,\quad m_{i,j}\in\mathbb{R},\quad\forall i=1,\dots,N_{\mathsf{u}},\quad\forall j=1,\dots,N_{\mathsf{p}},$ (A.4)
with $\mathrm{u}_{1}^{\beta_{1}}\,\mathrm{u}_{2}^{\beta_{2}}\,\cdots\,\mathrm{u}_{N_{\mathsf{u}}}^{\beta_{N_{\mathsf{u}}}}=1\Leftrightarrow\beta_{1}=\beta_{2}=\cdots=\beta_{N_{\mathsf{u}}}=0$, and all the exponents $m_{i,j}$ in (A.4) being explicitly known real numbers.
Aiming to minimize the number of involved quantities, one can also require the independence of the $N_{\mathsf{l}}$ coefficients $\lambda_{1},\dots,\lambda_{N_{\mathsf{l}}}$ (13), or, equivalently, verify that the rank of the matrix made by the exponents in (A.1) and (A.2) is equal to $N_{\mathsf{l}}$, i.e.,
$\mathrm{rank}\left(\begin{pmatrix}c_{1,1}&\cdots&c_{1,N_{\mathsf{x}}}&d_{1,1}&\cdots&d_{1,N_{\mathsf{p}}}\\ c_{2,1}&\cdots&c_{2,N_{\mathsf{x}}}&d_{2,1}&\cdots&d_{2,N_{\mathsf{p}}}\\ \vdots&\ddots&\vdots&\vdots&\ddots&\vdots\\ c_{N_{\mathsf{l}},1}&\cdots&c_{N_{\mathsf{l}},N_{\mathsf{x}}}&d_{N_{\mathsf{l}},1}&\cdots&d_{N_{\mathsf{l}},N_{\mathsf{p}}}\end{pmatrix}\right)=N_{\mathsf{l}}.$ (A.5)
In Section 5, we show that Properties A.1, A.2, A.4, A.5 and A.6 hold when considering the projectile model presented in Section 2.2. For such a physical model, we also assume Property A.3, as explicitly indicated by (28). Though such an assumption can be motivated by its dimensional consistency, there exist physical models satisfying Properties A.1, A.2, A.4, A.5 and A.6, while violating Property A.3 (see, for example, the equation for energy density (2.49) in [18]).
Appendix B Theorems
Theorem B.1 enables the computation of numerical values $\theta_{1},\dots,\theta_{N_{\mathsf{x}}}$ of the characteristic constants $\tilde{\theta}_{1},\dots,\tilde{\theta}_{N_{\mathsf{x}}}$ that minimize the deviation $C_{U}$ from unity of the magnitudes of $\lambda_{i}$ (13), with $i\in U\subseteq\{1,\dots,N_{\mathsf{l}}\}$.
Theorem B.1.
Given Property A.1 of coefficients $\lambda_{1},\dots,\lambda_{N_{\mathsf{l}}}$ (13), $U\subseteq\{1,\dots,N_{\mathsf{l}}\}$ and any fixed choice of $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$, the values of $(\theta_{1},\dots,\theta_{N_{\mathsf{x}}})\in(0,\infty)^{N_{\mathsf{x}}}$ attaining the global minimum of the function
$C_{U}(\theta_{1},\dots,\theta_{N_{\mathsf{x}}})\coloneqq\sum_{i\in U}\left(\log_{10}(\lambda_{i}(\tilde{\theta}_{1},\dots,\tilde{\theta}_{N_{\mathsf{x}}},\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}))\right)^{2},\quad C_{U}:(0,\infty)^{N_{\mathsf{x}}}\to[0,\infty),$ (B.1)
correspond to the solutions of the following linear system of $N_{\mathsf{x}}$ equations with $N_{\mathsf{x}}$ unknowns $\rho_{1},\dots,\rho_{N_{\mathsf{x}}}\in\mathbb{R}$:
$\sum_{k=1}^{N_{\mathsf{x}}}\left(\sum_{i\in U}\,c_{i,k}\,c_{i,j}\right)\rho_{k}=-\sum_{i\in U}\,c_{i,j}\,\log_{10}(\kappa_{i}),\quad\forall j=1,\dots,N_{\mathsf{x}},$ (B.2)
where $\rho_{k}\coloneqq\log_{10}(\theta_{k})$ and the numerical value $\kappa_{i}\in(0,\infty)$ of each $\tilde{\kappa}_{i}=\tilde{\kappa}_{i}(\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}})$ is fixed.
Proof.
By employing the notation (1) in (A.1) (Property A.1), one has
$\lambda_{i}=\tilde{\kappa}_{i}\,\tilde{\theta}_{1}^{c_{i,1}}\,\tilde{\theta}_{2}^{c_{i,2}}\,\cdots\,\tilde{\theta}_{N_{\mathsf{x}}}^{c_{i,N_{\mathsf{x}}}}=\kappa_{i}[\tilde{\kappa}_{i}]\,\theta_{1}^{c_{i,1}}[\tilde{\theta}_{1}]^{c_{i,1}}\,\theta_{2}^{c_{i,2}}[\tilde{\theta}_{2}]^{c_{i,2}}\,\cdots\,\theta_{N_{\mathsf{x}}}^{c_{i,{N_{\mathsf{x}}}}}[\tilde{\theta}_{N_{\mathsf{x}}}]^{c_{i,{N_{\mathsf{x}}}}},\quad\forall i=1,\dots,N_{\mathsf{l}}.$ (B.3)
As each of $\lambda_{1},\dots,\lambda_{N_{\mathsf{l}}}$ is dimensionless, i.e.,
$[\lambda_{i}]=[\tilde{\kappa}_{i}]\,[\tilde{\theta}_{1}]^{c_{i,1}}\,[\tilde{\theta}_{2}]^{c_{i,2}}\,\cdots\,[\tilde{\theta}_{N_{\mathsf{x}}}]^{c_{i,{N_{\mathsf{x}}}}}=1,\quad\forall i=1,\dots,N_{\mathsf{l}},$ (B.4)
it follows from (A.1) and (B.3) that
$\lambda_{i}=\kappa_{i}\,\theta_{1}^{c_{i,1}}\,\theta_{2}^{c_{i,2}}\,\cdots\,\theta_{N_{\mathsf{x}}}^{c_{i,{N_{\mathsf{x}}}}}\in(0,\infty),\quad\forall i=1,\dots,N_{\mathsf{l}}.$ (B.5)
Then, (B.5) allows writing the function $C_{U}$ as
$C_{U}(\vec{\rho})=\sum_{i\in U}\left(\log_{10}(\kappa_{i})+\sum_{k=1}^{N_{\mathsf{x}}}\,c_{i,k}\,\rho_{k}\right)^{2},\quad\forall\vec{\rho}\coloneqq(\rho_{1},\dots,\rho_{N_{\mathsf{x}}})\in\mathbb{R}^{N_{\mathsf{x}}},$ (B.6)
where $\rho_{k}\coloneqq\log_{10}(\theta_{k})$ for any $k=1,\dots,N_{\mathsf{x}}$. The twice-differentiable function $C_{U}(\vec{\rho})$ (B.6) is convex because its Hessian matrix is positive semi-definite $\forall\vec{\rho}\in\mathbb{R}^{N_{\mathsf{x}}}$. Then, as guaranteed by Theorem 2.5 in [19], the global minimizers of $C_{U}$ can be found by imposing $\partial_{\rho_{j}}C_{U}=0$, $\forall j=1,\dots,N_{\mathsf{x}}$, yielding to the linear system (B.2) of $N_{\mathsf{x}}$ equations with $N_{\mathsf{x}}$ unknowns $\rho_{1},\dots,\rho_{N_{\mathsf{x}}}\in\mathbb{R}$.
Theorem B.2 algebraically formalizes a criterion for dimensionlessness of physical parameters $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$.
Theorem B.2.
Given Property A.6 of physical parameters $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$, any quantity of the form
$\tilde{p}_{1}^{z_{1}}\,\tilde{p}_{2}^{z_{2}}\,\cdots\,\tilde{p}_{N_{\mathsf{p}}}^{z_{N_{\mathsf{p}}}},\quad\mbox{with}\quad z_{1},z_{2},\dots,z_{N_{\mathsf{p}}}\in\mathbb{R},$ (B.7)
is dimensionless if and only if the vector $\vec{z}\coloneqq(z_{1},z_{2},\dots,z_{N_{\mathsf{p}}})^{T}\in\mathbb{R}^{N_{\mathsf{p}}}$ of exponents in (B.7) belongs to the null space (kernel) of the matrix
$\textbf{M}\coloneqq\begin{pmatrix}m_{1,1}&m_{1,2}&\cdots&m_{1,N_{\mathsf{p}}}\\ m_{2,1}&m_{2,2}&\cdots&m_{2,N_{\mathsf{p}}}\\ \vdots&\vdots&\ddots&\vdots\\ m_{N_{\mathsf{u}},1}&m_{N_{\mathsf{u}},2}&\cdots&m_{N_{\mathsf{u}},N_{\mathsf{p}}}\end{pmatrix}\in\mathbb{R}^{N_{\mathsf{u}}\times N_{\mathsf{p}}},$ (B.8)
where the entries of M (B.8) are the exponents in (A.4) specified by Property A.6.
Proof.
Any quantity of the form (B.7) is unitless if and only if its dimensions are
$[\tilde{p}_{1}]^{z_{1}}\,[\tilde{p}_{2}]^{z_{2}}\,\cdots\,[\tilde{p}_{N_{\mathsf{p}}}]^{z_{N_{\mathsf{p}}}}=1,$ (B.9)
or, equivalently, by using (A.4) (Property A.6),
$[\tilde{p}_{1}]^{z_{1}}\,[\tilde{p}_{2}]^{z_{2}}\,\cdots\,[\tilde{p}_{N_{\mathsf{p}}}]^{z_{N_{\mathsf{p}}}}=\mathrm{u}_{1}^{\sum_{j=1}^{N_{\mathsf{p}}}m_{1,j}\,z_{j}}\,\mathrm{u}_{2}^{\sum_{j=1}^{N_{\mathsf{p}}}m_{2,j}\,z_{j}}\,\cdots\,\mathrm{u}_{N_{\mathsf{u}}}^{\sum_{j=1}^{N_{\mathsf{p}}}m_{N_{\mathsf{u}},j}\,z_{j}}=1.$ (B.10)
From independence of $\mathrm{u}_{1},\dots,\mathrm{u}_{N_{\mathsf{u}}}$ assured by Property A.6, (B.10) is equivalent to
$\sum_{j=1}^{N_{\mathsf{p}}}m_{i,j}\,z_{j}=0,\quad\forall i=1,\dots,N_{\mathsf{u}},\quad\mbox{i.e.,}\quad\textbf{M}\,\vec{z}=\vec{0},$ (B.11)
for M defined in (B.8) and $\vec{z}\coloneqq(z_{1},z_{2},\dots,z_{N_{\mathsf{p}}})^{T}\in\mathbb{R}^{N_{\mathsf{p}}}$. In other words, any quantity of the form (B.7) is dimensionless if and only if the vector $\vec{z}$ of exponents in (B.7) belongs to the null space (kernel) of matrix M (B.8), i.e., $\vec{z}\in\mathrm{ker}(\textbf{M})$.
Theorem B.3 determines the minimal number of independent dimensionless parameters and indicates a way to compute them in terms of the provided physical parameters $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$.
Theorem B.3.
Given Properties A.4 to A.6 of physical parameters $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$, there exist only $N_{\mathsf{0}}=N_{\mathsf{p}}-\mathrm{rank}(\textbf{M})$ independent dimensionless parameters $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}\in(0,\infty)$ of the form (B.7), with M as in (B.8). Such parameters $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}$ are identified by the vectors composing a basis of the null space (kernel) of the matrix M (B.8):
$\pi_{k}=\tilde{p}_{1}^{b_{k,1}}\,\tilde{p}_{2}^{b_{k,2}}\,\cdots\,\tilde{p}_{N_{\mathsf{p}}}^{b_{k,N_{\mathsf{p}}}},\quad\vec{b}_{k}\coloneqq(b_{k,1},b_{k,2},\dots,b_{k,N_{\mathsf{p}}})^{T}\in\mathbb{R}^{N_{\mathsf{p}}},$
$\mbox{with}\quad\vec{b}_{1},\vec{b}_{2},\dots,\vec{b}_{N_{\mathsf{0}}}\quad\mbox{constituing a}\quad\mathrm{basis}(\mathrm{ker}(\textbf{M})),$ (B.12)
and $b_{k,j}$ being the $j$-th entry of the vector $\vec{b}_{k}$ for any $k=1,\dots,N_{\mathsf{0}}$ and $j=1,\dots,N_{\mathsf{p}}$.
Proof.
First, we show that Properties A.4, A.5 and A.6 imply that the parameters $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}$ (B.12) are dimensionless, strictly positive and independent. Then, we conclude the proof of Theorem B.3 by demonstrating that any dimensionless quantity of the form (B.7) can be written in terms of such $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}$ only, as this means that there are no further independent quantities one can consider.
Given the matrix M (B.8) of exponents in (A.4) (Property A.6), any basis of $\mathrm{ker}(\textbf{M})$ is composed by $N_{\mathsf{0}}\coloneqq\mathrm{dim}(\mathrm{ker}(\textbf{M}))=N_{\mathsf{p}}-\mathrm{rank}(\textbf{M})$ independent vectors $\vec{b}_{1},\dots,\vec{b}_{N_{\mathsf{0}}}\in\mathrm{ker}(\textbf{M})\subseteq\mathbb{R}^{N_{\mathsf{p}}}$. The corresponding $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}$ (B.12) are dimensionless since $\vec{b}_{1},\dots,\vec{b}_{N_{\mathsf{0}}}\in\mathrm{ker}(\textbf{M})$, as guaranteed by Theorem B.2 assuming Property A.6.
Then, by employing the notation (1) in (B.12), one has
$\pi_{k}=\tilde{p}_{1}^{b_{k,1}}\,\cdots\,\tilde{p}_{N_{\mathsf{p}}}^{b_{k,N_{\mathsf{p}}}}=p_{1}^{b_{k,1}}\,[\tilde{p}_{1}]^{b_{k,1}}\,\cdots\,p_{N_{\mathsf{p}}}^{b_{k,{N_{\mathsf{p}}}}}\,[\tilde{p}_{N_{\mathsf{p}}}]^{b_{k,{N_{\mathsf{p}}}}},\quad\forall k=1,\dots,N_{\mathsf{0}}.$ (B.13)
As $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}$ are dimensionless, i.e.,
$[\pi_{k}]=[\tilde{p}_{1}]^{b_{k,1}}\,\cdots\,[\tilde{p}_{N_{\mathsf{p}}}]^{b_{k,{N_{\mathsf{p}}}}}=1,\quad\forall k=1,\dots,N_{\mathsf{0}},$ (B.14)
and $p_{1},\dots,p_{N_{\mathsf{p}}}\in(0,\infty)$ (Property A.5), it follows the strict positiveness of $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}$ (B.12), i.e.,
$\pi_{k}=p_{1}^{b_{k,1}}\,\cdots\,p_{N_{\mathsf{p}}}^{b_{k,N_{\mathsf{p}}}}\in(0,\infty),\quad\forall k=1,\dots,N_{\mathsf{0}}.$ (B.15)
Moreover, such $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}$ (B.12) are independent since
$\pi_{1}^{a_{1}}\,\pi_{2}^{a_{2}}\,\cdots\,\pi_{N_{\mathsf{0}}}^{a_{N_{\mathsf{0}}}}=1\Leftrightarrow a_{1}=a_{2}=\cdots=a_{N_{\mathsf{0}}}=0.$ (B.16)
On the one hand, $a_{1}=a_{2}=\cdots=a_{N_{\mathsf{0}}}=0$ trivially implies that $\pi_{1}^{a_{1}}\,\pi_{2}^{a_{2}}\,\cdots\,\pi_{N_{\mathsf{0}}}^{a_{N_{\mathsf{0}}}}=1$. On the other hand, given $\pi_{1}^{a_{1}}\,\pi_{2}^{a_{2}}\,\cdots\,\pi_{N_{\mathsf{0}}}^{a_{N_{\mathsf{0}}}}=1$, it follows from (B.12) that
$\pi_{1}^{a_{1}}\,\pi_{2}^{a_{2}}\,\cdots\,\pi_{N_{\mathsf{0}}}^{a_{N_{\mathsf{0}}}}=\tilde{p}_{1}^{\sum_{k=1}^{N_{\mathsf{0}}}a_{k}\,b_{k,1}}\,\tilde{p}_{2}^{\sum_{k=1}^{N_{\mathsf{0}}}a_{k}\,b_{k,2}}\,\cdots\,\tilde{p}_{N_{\mathsf{p}}}^{\sum_{k=1}^{N_{\mathsf{0}}}a_{k}\,b_{k,N_{\mathsf{p}}}}\,=1,$ (B.17)
which implies, from independence of $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$ assured by Property A.4,
$\sum_{k=1}^{N_{\mathsf{0}}}a_{k}\,b_{k,j}=0,\quad\forall j=1,\dots,N_{\mathsf{p}},\quad\mbox{i.e.,}\quad a_{1}\,\vec{b}_{1}\,+\,a_{2}\,\vec{b}_{2}\,+\,\cdots\,a_{N_{\mathsf{0}}}\,\vec{b}_{N_{\mathsf{0}}}=\vec{0}.$ (B.18)
As $\vec{b}_{1},\dots,\vec{b}_{N_{\mathsf{0}}}$ constitute a basis of $\mathrm{ker}(\textbf{M})$, they are linearly independent vectors. Thus, (B.18) gives $a_{1}=a_{2}=\cdots=a_{N_{\mathsf{0}}}=0$ and (B.16) holds, showing that $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}$ (B.12) are independent.
Finally, any dimensionless quantity of the form $\tilde{p}_{1}^{z_{1}}\,\tilde{p}_{2}^{z_{2}}\,\cdots\,\tilde{p}_{N_{\mathsf{p}}}^{z_{N_{\mathsf{p}}}}$, with $z_{1},z_{2},\dots,z_{N_{\mathsf{p}}}\in\mathbb{R}$, is such that $\vec{z}\coloneqq(z_{1},z_{2},\dots,z_{N_{\mathsf{p}}})^{T}\in\mathrm{ker}(\textbf{M})$, as guaranteed by Theorem B.2 assuming Property A.6. Then, the vector $\vec{z}$ can be written as a linear combination of the vectors $\vec{b}_{1},\dots,\vec{b}_{N_{\mathsf{0}}}$ composing a basis of $\mathrm{ker}(\textbf{M})$. By using (B.12), it follows that
$\tilde{p}_{1}^{z_{1}}\,\tilde{p}_{2}^{z_{2}}\,\cdots\,\tilde{p}_{N_{\mathsf{p}}}^{z_{N_{\mathsf{p}}}}=\tilde{p}_{1}^{\sum_{k=1}^{N_{\mathsf{0}}}a_{k}\,b_{k,1}}\,\tilde{p}_{2}^{\sum_{k=1}^{N_{\mathsf{0}}}a_{k}\,b_{k,2}}\,\cdots\,\tilde{p}_{N_{\mathsf{p}}}^{\sum_{k=1}^{N_{\mathsf{0}}}a_{k}\,b_{k,N_{\mathsf{p}}}}\,=\pi_{1}^{a_{1}}\,\pi_{2}^{a_{2}}\,\cdots\,\pi_{N_{\mathsf{0}}}^{a_{N_{\mathsf{0}}}},$ (B.19)
for some $a_{1},\dots,a_{N_{\mathsf{0}}}\in\mathbb{R}$. In conclusion, any dimensionless quantity of the form (B.7) can be written in terms of $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}$ (B.12) only. ∎
Theorem B.4 states that the dimensionless coefficients $\lambda_{1},\dots,\lambda_{N_{\mathsf{l}}}$ (13) can be written in terms of the independent dimensionless parameters $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}$ (B.12) only.
Theorem B.4.
Given Properties A.1 to A.6 and $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}$ (B.12), each of the dimensionless coefficients $\lambda_{1},\dots,\lambda_{N_{\mathsf{l}}}$ (13) can be written as a power-law monomial of the independent dimensionless parameters $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}$ only, i.e.,
$\lambda_{i}=\pi_{1}^{\alpha_{i,1}}\,\pi_{2}^{\alpha_{i,2}}\,\cdots\,\pi_{N_{\mathsf{0}}}^{\alpha_{i,N_{\mathsf{0}}}},\quad\alpha_{i,k}\in\mathbb{R},\quad\forall i=1,\dots,N_{\mathsf{l}},\quad\forall k=1,\dots,N_{\mathsf{0}},$ (B.20)
where the exponents $\alpha_{i,k}$ in (B.20) are unknowns of scaling procedures.
Proof.
As it does for $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}$ (B.12) (see the proof of Theorem B.3), Property A.5 ensures that the numerical values $\kappa_{i}$ and $\theta_{j}$ of all $\tilde{\kappa}_{i}$ in (A.2) and $\tilde{\theta}_{j}$ in (A.3) are strictly positive. Then, it is possible to plug (A.2) and (A.3) into (A.1) (see Properties A.1, A.2 and A.3). This uncovers that each of the dimensionless coefficients $\lambda_{1},\dots,\lambda_{N_{\mathsf{l}}}$ (13) can be written as a power-law monomial of the physical parameters $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$ only, i.e., in the form (B.7):
$\exists\,\vec{l}_{i}\coloneqq(l_{i,1},l_{i,2},\dots,l_{i,N_{\mathsf{p}}})^{T}\in\mathbb{R}^{N_{\mathsf{p}}}\quad\mbox{such that}\quad\lambda_{i}=\tilde{p}_{1}^{l_{i,1}}\,\tilde{p}_{2}^{l_{i,2}}\,\cdots\,\tilde{p}_{N_{\mathsf{p}}}^{l_{i,N_{\mathsf{p}}}},$ (B.21)
being $l_{i,j}$ the $j$-th component of the vector $\vec{l}_{i}$ for any $i=1,\dots,N_{\mathsf{l}}$ and $j=1,\dots,N_{\mathsf{p}}$.
Each of the coefficients $\lambda_{1},\dots,\lambda_{N_{\mathsf{l}}}$ in (B.21) is dimensionless and, thus, Theorem B.2 guarantees, under Property A.6, that each of the vectors $\vec{l}_{1},\dots,\vec{l}_{N_{\mathsf{l}}}$ of exponents in (B.21) belongs to the kernel of the matrix M (B.8), i.e., $\vec{l}_{1},\dots,\vec{l}_{N_{\mathsf{l}}}\in\mathrm{ker}(\textbf{M})$. If one considers the basis’ vectors $\vec{b}_{1},\dots,\vec{b}_{N_{\mathsf{0}}}$ (B.12) (Theorem B.3) assuming Properties A.4, A.5 and A.6, any vector $\vec{l}_{i}$ (B.21) belonging to $\mathrm{ker}(\textbf{M})$ can be written (in a unique way) as a linear combination of such $\vec{b}_{1},\dots,\vec{b}_{N_{\mathsf{0}}}$ (B.12). Then, it follows
[…]
where $\alpha_{i,k}$ is the $k$-th component of the vector $\vec{\alpha}_{i}$, with $i=1,\dots,N_{\mathsf{l}}$ and $k=1,\dots,N_{\mathsf{0}}$.
By plugging (B.22) into (B.21) and, then, using (B.12) (Theorem B.3) assuming Properties A.4, A.5 and A.6, it follows that each of the coefficients $\lambda_{1},\dots,\lambda_{N_{\mathsf{l}}}$ (13) can be written as a power-law monomial of the independent dimensionless parameters $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}$ (B.12) only, i.e.,
$\lambda_{i}=\tilde{p}_{1}^{l_{i,1}}\,\cdots\,\tilde{p}_{N_{\mathsf{p}}}^{l_{i,N_{\mathsf{p}}}}=\tilde{p}_{1}^{\sum_{k=1}^{N_{\mathsf{0}}}\alpha_{i,k}\,b_{k,1}}\,\cdots\,\tilde{p}_{N_{\mathsf{p}}}^{\sum_{k=1}^{N_{\mathsf{0}}}\alpha_{i,k}\,b_{k,N_{\mathsf{p}}}}=\pi_{1}^{\alpha_{i,1}}\,\cdots\,\pi_{N_{\mathsf{0}}}^{\alpha_{i,N_{\mathsf{0}}}},$ (B.23)
being $l_{i,j}$ and $b_{k,j}$ the $j$-th components of the vectors $\vec{l}_{i}$ and $\vec{b}_{k}$, respectively, for any $i=1,\dots,N_{\mathsf{l}}$, $j=1,\dots,N_{\mathsf{p}}$ and $k=1,\dots,N_{\mathsf{0}}$. We remark that the values of the exponents $\alpha_{i,k}$ in (B.23) are not explicitly known and, in this sense, we state that they are unknowns of scaling procedures. ∎
Theorem B.5 offers a way to obtain the unknown exponents $t_{i,j}$ in (A.3).
Theorem B.5.
Given Properties A.1 to A.6, the vector
$\vec{t}\coloneqq\begin{pmatrix}\vec{t}_{1}\\ \vec{t}_{2}\\ \vdots\\ \vec{t}_{N_{\mathsf{x}}}\end{pmatrix}\in\mathbb{R}^{N_{\mathsf{p}}\,N_{\mathsf{x}}},\quad\mbox{with}\quad\vec{t}_{j}\coloneqq\begin{pmatrix}t_{1,j}\\ t_{2,j}\\ \vdots\\ t_{N_{\mathsf{p}},j}\end{pmatrix}\in\mathbb{R}^{N_{\mathsf{p}}},$ (B.24)
of unknown exponents in (A.3) must satisfy the linear system
$\textbf{Y}\,\vec{t}=\vec{\gamma},\quad\textbf{Y}\in\mathbb{R}^{(N_{\mathsf{l}}\,N_{\mathsf{p}})\times(N_{\mathsf{p}}\,N_{\mathsf{x}})},\quad\vec{\gamma}\in\mathbb{R}^{N_{\mathsf{l}}\,N_{\mathsf{p}}},$ (B.25)
where the entries of Y are explicitly known real numbers, while the entries of $\vec{\gamma}$ are explicit affine functions of the exponents $\alpha_{i,k}$ in (B.20), as specified by (B.38).
Proof.
We consider the vectors $\vec{l}_{1},\dots,\vec{l}_{N_{\mathsf{l}}}\in\mathbb{R}^{N_{\mathsf{p}}}$ of exponents in (B.21). As explained in the proof of Theorem B.4, we have that Theorems B.2 and B.3 combined with Properties A.4, A.5 and A.6 ensure (B.22) for each vector $\vec{l}_{1},\dots,\vec{l}_{N_{\mathsf{l}}}$. This can be understood as follows. Let us denote B as the matrix whose columns are the vectors $\vec{b}_{1},\dots,\vec{b}_{N_{\mathsf{0}}}$ (B.12), i.e.,
$\textbf{B}\coloneqq\begin{pmatrix}\vec{b}_{1}&\vec{b}_{2}&\cdots&\vec{b}_{N_{\mathsf{0}}}\end{pmatrix}\in\mathbb{R}^{N_{\mathsf{p}}\times N_{\mathsf{0}}},\quad\mbox{with}\quad N_{\mathsf{0}}<N_{\mathsf{p}}.$ (B.26)
Each of the vectors $\vec{l}_{1},\dots,\vec{l}_{N_{\mathsf{l}}}$ in (B.22) belongs to the column space of such a matrix B (B.26), which is equivalent to the orthogonal complement of the null space (kernel) of $\textbf{B}^{T}$, i.e.,
$\vec{l}_{i}\in\mathrm{colsp}(\textbf{B})\equiv\left(\mathrm{ker}(\textbf{B}^{T})\right)^{\bot},\quad\forall i=1,\dots,N_{\mathsf{l}}.$ (B.27)
One has $N_{\mathsf{0}}=N_{\mathsf{p}}-\mathrm{rank}(\textbf{M})<N_{\mathsf{p}}$ in (B.26) owing to $\mathrm{rank}(\textbf{M})\geq 1$. The latter inequality follows from Property A.6 guaranteeing that M (B.8) cannot be a matrix whose entries are all equal to zero. Then, since $\mathrm{rank}(\textbf{B}^{T})=\mathrm{rank}(\textbf{B})=N_{\mathsf{0}}$ and $\mathrm{dim}\left(\mathrm{ker}(\textbf{B}^{T})\right)=N_{\mathsf{p}}-N_{\mathsf{0}}>0$, we can consider $N_{\mathsf{p}}-N_{\mathsf{0}}$ vectors $\vec{v}_{1},\dots,\vec{v}_{N_{\mathsf{p}}-N_{\mathsf{0}}}\in\mathbb{R}^{N_{\mathsf{p}}}$ to form a basis of $\mathrm{ker}(\textbf{B}^{T})$. The assertion (B.27) is equivalently stated as each of $\vec{l}_{1},\dots,\vec{l}_{N_{\mathsf{l}}}$ being perpendicular to each of the basis’ vectors $\vec{v}_{1},\dots,\vec{v}_{N_{\mathsf{p}}-N_{\mathsf{0}}}$, i.e.,
$\textbf{V}^{T}\,\vec{l}_{i}=\vec{0},\quad\textbf{V}\coloneqq\begin{pmatrix}\vec{v}_{1}&\vec{v}_{2}&\cdots&\vec{v}_{N_{\mathsf{p}}-N_{\mathsf{0}}}\end{pmatrix}\in\mathbb{R}^{N_{\mathsf{p}}\times(N_{\mathsf{p}}-N_{\mathsf{0}})},\quad\forall i=1,\dots,N_{\mathsf{l}}.$ (B.28)
Property A.5 guarantees that the numerical values $\kappa_{i}$ and $\theta_{j}$ of all $\tilde{\kappa}_{i}$ in (A.2) and $\tilde{\theta}_{j}$ in (A.3) are strictly positive. Then, one can plug (A.2) and (A.3) into (A.1) to get
$\lambda_{i}=\tilde{p}_{1}^{\sum_{j=1}^{N_{\mathsf{x}}}c_{i,j}\,t_{1,j}\,+\,d_{i,1}}\,\tilde{p}_{2}^{\sum_{j=1}^{N_{\mathsf{x}}}c_{i,j}\,t_{2,j}\,+\,d_{i,2}}\,\cdots\,\tilde{p}_{N_{\mathsf{p}}}^{\sum_{j=1}^{N_{\mathsf{x}}}c_{i,j}\,t_{N_{\mathsf{p}},j}\,+\,d_{i,N_{\mathsf{p}}}},\quad\forall i=1,\dots,N_{\mathsf{l}}.$ (B.29)
By using (B.29) and the independence of physical parameters $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$ (Property A.4), the vectors $\vec{l}_{1},\dots,\vec{l}_{N_{\mathsf{l}}}$ of exponents in (B.21) can be computed in terms of $\vec{t}_{1},\dots,\vec{t}_{N_{\mathsf{x}}}$ (B.24):
$\vec{l}_{i}=c_{i,1}\,\vec{t}_{1}\,+\,c_{i,2}\,\vec{t}_{2}\,+\,\cdots\,+\,c_{i,N_{\mathsf{x}}}\,\vec{t}_{N_{\mathsf{x}}}\,+\,\vec{d}_{i},\quad\vec{d}_{i}\coloneqq(d_{i,1},d_{i,2},\dots,d_{i,{N_{\mathsf{p}}}})^{T},\quad\forall i=1,\dots,N_{\mathsf{l}}.$ (B.30)
Inserting $\vec{l}_{i}$ (B.30) into (B.28) gives
$c_{i,1}\,\textbf{V}^{T}\,\vec{t}_{1}\,+\,c_{i,2}\,\textbf{V}^{T}\,\vec{t}_{2}\,+\,\cdots\,+\,c_{i,N_{\mathsf{x}}}\,\textbf{V}^{T}\,\vec{t}_{N_{\mathsf{x}}}=-\textbf{V}^{T}\,\vec{d}_{i},\quad\forall i=1,\dots,N_{\mathsf{l}},$ (B.31)
that reads in terms of the vector $\vec{t}$ (B.24) of unknowns as
$\textbf{T}\,\vec{t}=\vec{\tau},\quad\textbf{T}\in\mathbb{R}^{N_{\mathsf{t}}\times(N_{\mathsf{p}}\,N_{\mathsf{x}})},\quad\vec{\tau}\in\mathbb{R}^{N_{\mathsf{t}}},\quad N_{\mathsf{t}}\coloneqq N_{\mathsf{l}}\,(N_{\mathsf{p}}-N_{\mathsf{0}}),$ (B.32)
$\textbf{T}=\begin{pmatrix}c_{1,1}\,\textbf{V}^{T}&c_{1,2}\,\textbf{V}^{T}&\cdots&c_{1,N_{\mathsf{x}}}\,\textbf{V}^{T}\\ c_{2,1}\,\textbf{V}^{T}&c_{2,2}\,\textbf{V}^{T}&\cdots&c_{2,N_{\mathsf{x}}}\,\textbf{V}^{T}\\ \vdots&\vdots&\ddots&\vdots\\ c_{N_{\mathsf{l}},1}\,\textbf{V}^{T}&c_{N_{\mathsf{l}},2}\,\textbf{V}^{T}&\cdots&c_{N_{\mathsf{l}},N_{\mathsf{x}}}\,\textbf{V}^{T}\end{pmatrix},\quad\vec{\tau}=-\begin{pmatrix}\textbf{V}^{T}\,\vec{d}_{1}\\ \textbf{V}^{T}\,\vec{d}_{2}\\ \vdots\\ \textbf{V}^{T}\,\vec{d}_{N_{\mathsf{l}}}\end{pmatrix},$ (B.33)
with V and $\vec{d}_{1},\dots,\vec{d}_{N_{\mathsf{l}}}$ defined as in (B.28) and (B.30), respectively. By using any solution $\vec{t}\coloneqq(\vec{t}_{1},\vec{t}_{2},\dots,\vec{t}_{N_{\mathsf{x}}})^{T}$ to (B.32)-(B.33) in (B.30), the resulting vectors $\vec{l}_{1},\dots,\vec{l}_{N_{\mathsf{l}}}$ (B.30) fulfill (B.28) that implies (B.27). In other words, such $\vec{l}_{1},\dots,\vec{l}_{N_{\mathsf{l}}}$ belong to the column space of the matrix B (B.26) generated by the basis’ vectors $\vec{b}_{1},\dots,\vec{b}_{N_{\mathsf{0}}}$ (B.12). Thus, each of the computed $\vec{l}_{1},\dots,\vec{l}_{N_{\mathsf{l}}}$ admits the unique vector $\vec{\alpha}_{i}$ (B.22) of coordinates with respect to the basis $\vec{b}_{1},\dots,\vec{b}_{N_{\mathsf{0}}}$ (B.12). That is to say, each of the $N_{\mathsf{l}}$ linear systems written in (B.22), i.e., $\textbf{B}\,\vec{\alpha}_{i}=\vec{l}_{i}$, with $i=1,\dots,N_{\mathsf{l}}$, can be solved providing the unique solution $\vec{\alpha}_{i}\in\mathbb{R}^{N_{\mathsf{0}}}$:
$\vec{\alpha}_{i}=\textbf{B}^{+}\,\vec{l}_{i},\quad\textbf{B}^{+}\coloneqq\left(\textbf{B}^{T}\,\textbf{B}\,\right)^{-1}\textbf{B}^{T}\in\mathbb{R}^{N_{\mathsf{0}}\times N_{\mathsf{p}}},\quad\forall i=1,\dots,N_{\mathsf{l}},$ (B.34)
where $\textbf{B}^{+}$ is the pseudo-inverse of the rectangular matrix B (B.26). The inverse matrix $\left(\textbf{B}^{T}\,\textbf{B}\,\right)^{-1}\in\mathbb{R}^{N_{\mathsf{0}}\times N_{\mathsf{0}}}$ in (B.34) is well defined because $\mathrm{rank}(\textbf{B})=N_{\mathsf{0}}<N_{\mathsf{p}}$ implies that $\mathrm{rank}(\textbf{B}^{T}\,\textbf{B})=N_{\mathsf{0}}$. By plugging $\vec{l}_{i}$ (B.30) into (B.34), it follows
$c_{i,1}\,\textbf{B}^{+}\,\vec{t}_{1}\,+\,c_{i,2}\,\textbf{B}^{+}\,\vec{t}_{2}\,+\,\cdots\,+\,c_{i,N_{\mathsf{x}}}\,\textbf{B}^{+}\,\vec{t}_{N_{\mathsf{x}}}=\vec{\alpha}_{i}\,-\,\textbf{B}^{+}\,\vec{d}_{i},\quad\forall i=1,\dots,N_{\mathsf{l}},$ (B.35)
that reads in terms of the vector $\vec{t}$ (B.24) of unknowns as
$\textbf{A}\,\vec{t}=\vec{\alpha}\,-\,\vec{\omega},\quad\textbf{A}\in\mathbb{R}^{(N_{\mathsf{l}}\,N_{\mathsf{0}})\times(N_{\mathsf{p}}\,N_{\mathsf{x}})},\quad\vec{\alpha},\vec{\omega}\in\mathbb{R}^{N_{\mathsf{l}}\,N_{\mathsf{0}}},$ (B.36)
$\textbf{A}=\begin{pmatrix}c_{1,1}\,\textbf{B}^{+}&c_{1,2}\,\textbf{B}^{+}&\cdots&c_{1,N_{\mathsf{x}}}\,\textbf{B}^{+}\\ c_{2,1}\,\textbf{B}^{+}&c_{2,2}\,\textbf{B}^{+}&\cdots&c_{2,N_{\mathsf{x}}}\,\textbf{B}^{+}\\ \vdots&\vdots&\ddots&\vdots\\ c_{N_{\mathsf{l}},1}\,\textbf{B}^{+}&c_{N_{\mathsf{l}},2}\,\textbf{B}^{+}&\cdots&c_{N_{\mathsf{l}},N_{\mathsf{x}}}\,\textbf{B}^{+}\end{pmatrix},\quad\vec{\omega}=\begin{pmatrix}\textbf{B}^{+}\,\vec{d}_{1}\\ \textbf{B}^{+}\,\vec{d}_{2}\\ \vdots\\ \textbf{B}^{+}\,\vec{d}_{N_{\mathsf{l}}}\end{pmatrix},$ (B.37)
where $\vec{\alpha}$ is defined as in (B.39), being the vector $\in\mathbb{R}^{N_{\mathsf{l}}\,N_{\mathsf{0}}}$ whose entries are the exponents $\alpha_{i,k}$ in (B.20). The matrix $\textbf{B}^{+}$ and the vectors $\vec{d}_{1},\dots,\vec{d}_{N_{\mathsf{l}}}$ are defined as in (B.34) and (B.30), respectively. Both linear systems (B.32)-(B.33) and (B.36)-(B.37) must be satisfied by the vector $\vec{t}$ (B.24) of unknowns, yielding to the matrix equation $\textbf{Y}\,\vec{t}=\vec{\gamma}$ (B.25) for $\vec{t}$, with Y and $\vec{\gamma}$ given by
$\textbf{Y}=\begin{pmatrix}\textbf{A}\\ \textbf{T}\end{pmatrix}\in\mathbb{R}^{(N_{\mathsf{l}}\,N_{\mathsf{p}})\times(N_{\mathsf{p}}\,N_{\mathsf{x}})},\quad\vec{\gamma}=\begin{pmatrix}\vec{\alpha}-\vec{\omega}\\ \vec{\tau}\end{pmatrix}\in\mathbb{R}^{N_{\mathsf{l}}\,N_{\mathsf{p}}}.$ (B.38)
The entries of Y (B.38) are explicitly known real numbers, as given by those of A (B.37) and T (B.33). Since the components of $\vec{\omega}$ (B.37) and $\vec{\tau}$ (B.33) are explicitly known, the entries of $\vec{\gamma}$ (B.38) are explicit affine functions of the exponents $\alpha_{i,k}$ in (B.20), collected by the vector $\vec{\alpha}$ (B.39).
Theorem B.6 provides a way for finding unknown exponents $\alpha_{i,k}$ in (B.20) (Theorem B.4).
Theorem B.6.
Given Properties A.1 to A.6 and the consequent expression (B.20) of coefficients $\lambda_{1},\dots,\lambda_{N_{\mathsf{l}}}$ (13), the vector
$\vec{\alpha}\coloneqq\begin{pmatrix}\vec{\alpha}_{1}\\ \vec{\alpha}_{2}\\ \vdots\\ \vec{\alpha}_{N_{\mathsf{l}}}\end{pmatrix}\in\mathbb{R}^{N_{\mathsf{l}}\,N_{\mathsf{0}}},\quad\mbox{with}\quad\vec{\alpha}_{i}\coloneqq\begin{pmatrix}\alpha_{i,1}\\ \alpha_{i,2}\\ \vdots\\ \alpha_{i,N_{\mathsf{0}}}\end{pmatrix}\in\mathbb{R}^{N_{\mathsf{0}}},$ (B.39)
of exponents in (B.20) must satisfy the linear system
$\textbf{S}\,\vec{\alpha}=\vec{s},\quad\textbf{S}\in\mathbb{R}^{N_{\mathsf{r}}\times N_{\mathsf{c}}},\quad\vec{s}\in\mathbb{R}^{N_{\mathsf{r}}},\quad N_{\mathsf{c}}\coloneqq N_{\mathsf{l}}\,N_{\mathsf{0}},\quad N_{\mathsf{r}}\coloneqq N_{\mathsf{l}}\,N_{\mathsf{p}}-\mathrm{rank}(\textbf{Y}),$ (B.40)
where the entries of S and $\vec{s}$ are explicitly known real numbers identified in (B.45)-(B.46) and the components of the matrix $\textbf{Y}\in\mathbb{R}^{(N_{\mathsf{l}}\,N_{\mathsf{p}})\times(N_{\mathsf{p}}\,N_{\mathsf{x}})}$ are explicitly known, as specified by (B.38).
Proof.
Assuming Property A.3, there exists at least a vector $\vec{t}\coloneqq\left(t_{1,1},\dots,t_{N_{\mathsf{p}},N_{\mathsf{x}}}\right)^{T}\in\mathbb{R}^{N_{\mathsf{p}}\,N_{\mathsf{x}}}$ of exponents that ensures (A.3). Under Properties A.1, A.2, A.3, A.4, A.5 and A.6, Theorem B.5 guarantees that such a vector $\vec{t}$ (B.24) satisfies the linear system $\textbf{Y}\,\vec{t}=\vec{\gamma}$ (B.25). Then, there must be at least a solution to such a matrix equation (B.25). Equivalently, the vector $\vec{\gamma}\in\mathbb{R}^{N_{\mathsf{l}}\,N_{\mathsf{p}}}$ must belong to the column space of the matrix $\textbf{Y}\in\mathbb{R}^{(N_{\mathsf{l}}\,N_{\mathsf{p}})\times(N_{\mathsf{p}}\,N_{\mathsf{x}})}$, which is the orthogonal complement of the null space (kernel) of $\textbf{Y}^{T}$, i.e.,
$\vec{\gamma}\in\mathrm{colsp}(\textbf{Y})\equiv\left(\mathrm{ker}(\textbf{Y}^{T})\right)^{\bot}.$ (B.41)
Let us consider $N_{\mathsf{r}}\coloneqq N_{\mathsf{l}}\,N_{\mathsf{p}}-\mathrm{rank}(\textbf{Y})=\mathrm{dim}\left(\mathrm{ker}(\textbf{Y}^{T})\right)$ vectors $\vec{w}_{1},\dots,\vec{w}_{N_{\mathsf{r}}}\in\mathbb{R}^{N_{\mathsf{l}}\,N_{\mathsf{p}}}$ to form a basis of the kernel of $\textbf{Y}^{T}$ and the corresponding matrix
$\textbf{W}\coloneqq\begin{pmatrix}\vec{w}_{1}&\vec{w}_{2}&\cdots&\vec{w}_{N_{\mathsf{r}}}\end{pmatrix}\in\mathbb{R}^{(N_{\mathsf{l}}\,N_{\mathsf{p}})\times N_{\mathsf{r}}}.$ (B.42)
For the matrix W (B.42), the statement (B.41) is equivalent to
$\textbf{W}^{T}\,\vec{\gamma}=\vec{0},$ (B.43)
or, given $\vec{\gamma}$ as in (B.38) and $w_{i,j}$ as the entry of row $i$ and column $j$ of $\textbf{W}^{T}\in\mathbb{R}^{N_{\mathsf{r}}\times(N_{\mathsf{l}}\,N_{\mathsf{p}})}$, to
$\begin{pmatrix}w_{1,1}&w_{1,2}&\cdots&w_{1,N_{\mathsf{l}}N_{\mathsf{p}}}\\ w_{2,1}&w_{2,2}&\cdots&w_{2,N_{\mathsf{l}}N_{\mathsf{p}}}\\ \vdots&\vdots&\ddots&\vdots\\ w_{N_{\mathsf{r}},1}&w_{N_{\mathsf{r}},2}&\cdots&w_{N_{\mathsf{r}},N_{\mathsf{l}}N_{\mathsf{p}}}\end{pmatrix}\begin{pmatrix}\vec{\alpha}-\vec{\omega}\\ \vec{\tau}\end{pmatrix}=\vec{0},$ (B.44)
where $\vec{\alpha}\in\mathbb{R}^{N_{\mathsf{l}}\,N_{\mathsf{0}}}$ (B.39) is the vector of exponents in (B.20), while $\vec{\omega}\in\mathbb{R}^{N_{\mathsf{l}}\,N_{\mathsf{0}}}$ and $\vec{\tau}\in\mathbb{R}^{N_{\mathsf{l}}\,(N_{\mathsf{p}}-N_{\mathsf{0}})}$ are explicitly known, as specified by (B.37) and (B.33), respectively. By using $N_{\mathsf{0}}=N_{\mathsf{p}}-\mathrm{rank}(\textbf{M})$ from Theorem B.3 under Properties A.4, A.5 and A.6, we remark that $N_{\mathsf{c}}\coloneqq N_{\mathsf{l}}\,N_{\mathsf{0}}=N_{\mathsf{l}}\,(N_{\mathsf{p}}-\mathrm{rank}(\textbf{M}))<N_{\mathsf{l}}\,N_{\mathsf{p}}$, being M (B.8) defined by Theorem B.2 under Property A.6. The latter inequality follows from Property A.6 guaranteeing that M (B.8) cannot be a matrix whose entries are all equal to zero, i.e., $\mathrm{rank}(\textbf{M})\geq 1$. Then, it is possible to write (B.44) in terms of $\vec{\alpha}\in\mathbb{R}^{N_{\mathsf{c}}}$ as the matrix equation $\textbf{S}\,\vec{\alpha}=\vec{s}$ (B.40), where the matrix S and the vector $\vec{s}$ are
$\textbf{S}=\begin{pmatrix}w_{1,1}&w_{1,2}&\cdots&w_{1,N_{\mathsf{l}}N_{\mathsf{0}}}\\ w_{2,1}&w_{2,2}&\cdots&w_{2,N_{\mathsf{l}}N_{\mathsf{0}}}\\ \vdots&\vdots&\ddots&\vdots\\ w_{N_{\mathsf{r}},1}&w_{N_{\mathsf{r}},2}&\cdots&w_{N_{\mathsf{r}},N_{\mathsf{l}}N_{\mathsf{0}}}\end{pmatrix}\in\mathbb{R}^{N_{\mathsf{r}}\times(N_{\mathsf{l}}\,N_{\mathsf{0}})},$ (B.45)
$\vec{s}=\begin{pmatrix}w_{1,1}&w_{1,2}&\cdots&w_{1,N_{\mathsf{l}}N_{\mathsf{0}}}\\ w_{2,1}&w_{2,2}&\cdots&w_{2,N_{\mathsf{l}}N_{\mathsf{0}}}\\ \vdots&\vdots&\ddots&\vdots\\ w_{N_{\mathsf{r}},1}&w_{N_{\mathsf{r}},2}&\cdots&w_{N_{\mathsf{r}},N_{\mathsf{l}}N_{\mathsf{0}}}\end{pmatrix}\vec{\omega}-\begin{pmatrix}w_{1,N_{\mathsf{l}}N_{\mathsf{0}}+1}&w_{1,N_{\mathsf{l}}N_{\mathsf{0}}+2}&\cdots&w_{1,N_{\mathsf{l}}N_{\mathsf{p}}}\\ w_{2,N_{\mathsf{l}}N_{\mathsf{0}}+1}&w_{2,N_{\mathsf{l}}N_{\mathsf{0}}+2}&\cdots&w_{2,N_{\mathsf{l}}N_{\mathsf{p}}}\\ \vdots&\vdots&\ddots&\vdots\\ w_{N_{\mathsf{r}},N_{\mathsf{l}}N_{\mathsf{0}}+1}&w_{N_{\mathsf{r}},N_{\mathsf{l}}N_{\mathsf{0}}+2}&\cdots&w_{N_{\mathsf{r}},N_{\mathsf{l}}N_{\mathsf{p}}}\end{pmatrix}\vec{\tau}\in\mathbb{R}^{N_{\mathsf{r}}},$ (B.46)
with $w_{i,j}$ being the explicitly known entry of row $i$ and column $j$ of the transpose of the matrix W (B.42), while $\vec{\omega}\in\mathbb{R}^{N_{\mathsf{l}}\,N_{\mathsf{0}}}$ and $\vec{\tau}\in\mathbb{R}^{N_{\mathsf{l}}\,(N_{\mathsf{p}}-N_{\mathsf{0}})}$ being explicitly given by (B.37) and (B.33), respectively.
Theorem B.7 enables the computation of numerical values of the exponents $\alpha_{i,k}$ in (B.20) that minimize the deviation $C_{U}$ from unity of the magnitudes of $\lambda_{i}$, with $i\in U\subseteq\{1,\dots,N_{\mathsf{l}}\}$.
Theorem B.7.
Given any fixed choice of $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}\in(0,\infty)$ and set $U\subseteq\{1,\dots,N_{\mathsf{l}}\}$, the values of $\vec{\alpha}\in\mathbb{R}^{N_{\mathsf{c}}}$ (B.39), $N_{\mathsf{c}}\coloneqq N_{\mathsf{l}}\,N_{\mathsf{0}}$, attaining the global minimum of the function
$C_{U}(\vec{\alpha})\coloneqq\sum_{i\in U}\left(\log_{10}(\lambda_{i})\right)^{2},\quad\lambda_{i}=\pi_{1}^{\alpha_{i,1}}\,\pi_{2}^{\alpha_{i,2}}\,\cdots\,\pi_{N_{\mathsf{0}}}^{\alpha_{i,N_{\mathsf{0}}}},\quad C_{U}:\mathbb{R}^{N_{\mathsf{c}}}\to[0,\infty),$ (B.47)
subject to a constraint of the form
$\boldsymbol{\Sigma}\,\vec{\alpha}=\vec{\sigma},\quad\boldsymbol{\Sigma}\in\mathbb{R}^{\hat{N_{\mathsf{r}}}\times N_{\mathsf{c}}},\quad\vec{\sigma}\in\mathbb{R}^{\hat{N_{\mathsf{r}}}},\quad\hat{N_{\mathsf{r}}}\in\mathbb{N},$ (B.48)
satisfy the system of equations
$\boldsymbol{\Sigma}\,\vec{\alpha}=\vec{\sigma},\quad\nabla C_{U}(\vec{\alpha})\,+\,\boldsymbol{\Sigma}^{T}\,\vec{\nu}=\vec{0},$ (B.49)
assuming that
$\mathrm{rank}(\boldsymbol{\Sigma})=\hat{N_{\mathsf{r}}}<N_{\mathsf{c}},$ (B.50)
and there is a vector $\vec{\nu}\in\mathbb{R}^{\hat{N_{\mathsf{r}}}}$ such that (B.49) holds.
Proof.
The twice continuously differentiable function $C_{U}=C_{U}(\vec{\alpha})$ in (B.47) is convex as its Hessian matrix is diagonal with non-negative entries $\forall\vec{\alpha}\in\mathbb{R}^{N_{\mathsf{c}}}$. Suppose one also aims to satisfy a constraint of the form (B.48). It follows that finding any minimum of $C_{U}=C_{U}(\vec{\alpha})$ reads as a convex optimization problem with equality constraints, as considered in Section 10.1 of [20].
Then, as guaranteed in Section 10.1 of [20] by assuming (B.50), the vector $\vec{\alpha}\in\mathbb{R}^{N_{\mathsf{c}}}$ is optimal for the problem under consideration if and only if there is a vector $\vec{\nu}\in\mathbb{R}^{\hat{N_{\mathsf{r}}}}$ such that (B.49) holds. Any found local minimizer is automatically a global minimizer, because the considered problem is convex, as discussed in Section 4.2.2 of [20].